### v4

# On the module of derivations of a line arrangement

Abstract.
To each multiple point $p$ in a line arrangement ${\mathcal{A}}$ in the complex projective plane we associate a derivation $\tilde{D}_{p}\in D_{0}({\mathcal{A}})$. We show first that these derivations span a large subspace of $D_{0}({\mathcal{A}})$, see Theorem 1.4 and especially Theorem 1.6. To each such derivation $\tilde{D}_{p}\in D_{0}({\mathcal{A}})$ we associate a polynomial $g_{p}$ which seems to play a key role in the characterization of the freeness of ${\mathcal{A}}$, see Theorem 1.8, as well as in the study of the position of the multiple points of ${\mathcal{A}}$ with respect to unions of lines, see Corollary 1.11 and Conjecture 1.12.
Key words and phrases:
line arrangement, Jacobian syzygies, derivations, combinatorics
2020 Mathematics Subject Classification:
Primary 14J70; Secondary 14B05, 32S05
1 partial support from the project “Singularities and Applications” - CF 132/31.07.2023 funded by the European Union - NextGenerationEU - through Romania’s National Recovery and Resilience Plan.
1. Introduction
Let $S=\mathbb{C}[x,y,z]$ be the graded polynomial ring in the variables $x,y,z$ with complex coefficients and let $f\in S_{d}$ be a homogeneous polynomial of degree $d$. Let $f_{x},f_{y}$ and $f_{z}$ denote the partial derivatives of $f$ with respect to $x,y$ and $z$ respectively. Recall that a Jacobian syzygy for $f$ is a triple
(1.1) $\rho=(a,b,c)\in S^{3}\text{ such that }af_{x}+bf_{y}+cf_{z}=0.$
With such a syzygy $\rho$, one can associate the following derivation
(1.2) $D_{\rho}=a{\partial_{x}}+b\partial_{y}+c\partial_{z}\in Der(S),$
of the polynomial ring $S$, such that $D_{\rho}(f)=0$, i.e. the derivation $D_{\rho}$ kills the polynomial $f$. Consider the graded $S$-module of Jacobian syzygies of $f$, namely
$Syz(f)=\{\rho=(a,b,c)\in S^{3}\text{ such that }af_{x}+bf_{y}+cf_{z}=0.\}.$
Note that $Syz(f)=D_{0}(f)$, the $S$-module of all derivations killing $f$, as defined in (1.2). In the sequel we often identify the syzygy $\rho$ with the corresponding derivation $D_{\rho}$. We say that the curve $C:f=0$ is an $m$-syzygy curve if the module $D_{0}(f)$ is minimally generated by $m$ homogeneous syzygies, say $\rho_{1},\rho_{2},...,\rho_{m}$, of degrees $d_{j}=\deg\rho_{j}$ ordered such that
$d_{1}\leq d_{2}\leq\ldots\leq d_{m}.$
We call these degrees $(d_{1},\ldots,d_{m})$ the exponents of the curve $C$. The smallest degree $d_{1}$ is sometimes denoted by ${\rm mdr}(f)$ and is called the minimal degree of a Jacobian relation for $f$.
In this paper we consider the case when $C$ is a line arrangement ${\mathcal{A}}:f=0$, and try to detect properties of the module $D_{0}(f)$ of derivations killing $f$ which depend only on the combinatorics of the arrangement ${\mathcal{A}}$. Our first main result is the following.
Theorem 1.1.
For the arrangements ${\mathcal{A}}:f=0$ of $d$ lines in $\mathbb{P}^{2}$ one has the following.
  1. (1)
    The dimension $\dim D_{0}(f)_{j}$ is combinatorially determined for any $j\geq d-3$. More precisely, for $j\geq d-3$, one has
    $\dim D_{0}(f)_{j}=\tau({\mathcal{A}})+3\binom{j+2}{2}-\binom{d+j+1}{2}.$
    In particular
    $\dim D_{0}(f)_{d-3}=\tau({\mathcal{A}})-\binom{d}{2}\text{ and }\dim D_{0}(f)_{d-2}=\tau({\mathcal{A}})-\binom{d-1}{2}.$
  2. (2)
    The dimension $\dim D_{0}(f)_{d-4}$ is not combinatorially determined in general.
As a direct consequence of Theorem 1.1 we have the following.
Corollary 1.2.
For any arrangement ${\mathcal{A}}:f=0$ of $d$ lines in $\mathbb{P}^{2}$ one has the following, where $m_{p}$ is the multiplicity of the point $p\in{\mathcal{A}}$ and the sum is over all points of multiplicity at least $3$.
  1. (1)
    $\dim D_{0}(f)_{d-3}=\sum_{p}\binom{m_{p}-1}{2}.$
    In particular $\dim D_{0}(f)_{d-3}=0$ if and only if ${\mathcal{A}}$ has only double points.
  2. (2)
    $\dim D_{0}(f)_{d-2}=\sum_{p}\binom{m_{p}-1}{2}+d-1.$
    In particular $\dim D_{0}(f)_{d-2}=d-1$ if and only if ${\mathcal{A}}$ has only double points.
A second theme of this paper is to construct some derivations in $D_{0}({\mathcal{A}})$ starting from the combinatorics of ${\mathcal{A}}$. More precisely, let ${\mathcal{A}}:f=0$ be a line arrangement of $d$ lines in $\mathbb{P}^{2}$ and let
$p=(u:v:w)$
be a point in ${\mathcal{A}}$ of multiplicity $m_{p}\geq 2$. We assume in this paper that ${\mathcal{A}}$ is not a pencil of lines, that is we assume that $m_{p}<d$. Then the defining equation $f=0$ can be written as
$f=f_{1p}f_{2p}=0,$
where $f_{1p}=0$ (resp. $f_{2p}=0$) is the equation of the subarrangement ${\mathcal{A}}_{1}$ (resp. ${\mathcal{A}}_{2}$) of ${\mathcal{A}}$ formed by all the lines in ${\mathcal{A}}$ passing (resp. not passing) through the point $p$. The following result tells that to the multiple point $p$ one can associate in a rather natural way a derivation in $D_{0}({\mathcal{A}})_{d-m_{p}}$. If $p$ is any point in the complex projective plane $\mathbb{P}^{2}$, and $D_{\rho}$ is a derivation as in (1.2), then we set
(1.3) $D_{\rho}(p)=a(p)\partial_{x}+b(p)\partial_{y}+c(p)\partial_{z},$
that is we evaluate all the coefficients of $D_{\rho}$ at $p$, and get a constant derivation, well defined up-to a nonzero constant factor.
Theorem 1.3.
For any arrangement ${\mathcal{A}}:f=0$ of $d$ lines in $\mathbb{P}^{2}$, and with the above notation, we consider the following two derivations
$D_{p}=u\partial_{x}+v\partial_{y}+w\partial_{z}$
and
$\tilde{D}_{p}=f_{2p}D_{p}-\frac{D_{p}(f_{2p})}{d}E,$
where $E=x{\partial_{x}}+y{\partial_{y}}+z{\partial_{z}}$ is the Euler derivation. Then the following hold.
  1. (1)
    $D_{p}\in D_{0}(f_{1p})$ and $\deg D_{p}=0$.
  2. (2)
    $\tilde{D}_{p}\in D_{0}(f)$ and $\deg\tilde{D}_{p}=d-m_{p}$.
  3. (3)
    $\tilde{D}_{p}(p)\neq 0$.
  4. (4)
    Let $q\neq p$ be another multiple point of ${\mathcal{A}}$. Then the coefficients of $\tilde{D}_{p}$ vanish at $q$ of order at least $m_{q}-2$.
  5. (5)
    For any point $q\in\mathbb{P}^{2}$, one has $\tilde{D}_{p}(q)=0$ if and only if $q$ is a multiple point of the arrangement ${\mathcal{A}}_{2}:f_{2p}=0$.
Such local derivations $\tilde{D}_{p}$ were already considered in our papers [7, proof of Theorem 1.2] and [8, proof of Lemma 3.2], but now we consider them in more details. The vector spaces spanned by the multiples of such local derivations is quite large in $D_{0}(f)$. For instance, we have the following result.
Theorem 1.4.
For any arrangement ${\mathcal{A}}:f=0$ of $d$ lines the vector space $D_{0}(f)_{d-3}$ is the direct sum
$\oplus_{p}S_{m_{p}-3}\tilde{D}_{p},$
where $p$ runs through the set of points $p\in{\mathcal{A}}$ of multiplicity $m_{p}\geq 3$.
The following result is a direct consequence of Theorem 1.4, to be compared with Theorem 1.1 $(2)$.
Corollary 1.5.
For any arrangement ${\mathcal{A}}:f=0$ of $d$ lines in $\mathbb{P}^{2}$ and for any $j\geq 4$, one has the following inequality
$\dim D_{0}(f)_{d-j}\geq\sum_{p}\binom{m_{p}-j+2}{2},$
where the sum is over all multiple points $p\in{\mathcal{A}}$ such that $m_{p}\geq j$.
The example of arrangement in the proof of Theorem 1.1 $(2)$ shows that the inequality in Corollary 1.5 may be strict. In relation to Theorem 1.4, we can prove the following key result.
Theorem 1.6.
For any arrangement ${\mathcal{A}}$ of $d$ lines in $\mathbb{P}^{2}$ which is not a pencil, and for any $j\geq d-2$, the local derivations $\tilde{D}_{p}$ generate $D_{0}({\mathcal{A}})_{j}$, when $p$ runs through all the multiple points of ${\mathcal{A}}$ of multiplicity $m_{p}\geq 2$. More precisely, one has
$D_{0}({\mathcal{A}})_{j}=\sum_{p}S_{j+m_{p}-d}\tilde{D}_{p},$
for any $j\geq d-2$, where this sum is not a direct sum in general.
Remark 1.7.
Since all the generators of the graded $S$-module have degrees $\leq d-2$, see for instance [22], it follows that to prove Theorem 1.6 it is enough to prove the case $j=d-2$.
Let ${\mathcal{A}}:f=0$ be a line arrangement and fix a minimal degree non-zero syzygy $\rho\in D_{0}(f)$, say $\rho=(a,b,c)$. For any homogeneous syzygy $\rho^{\prime}=(a^{\prime},b^{\prime},c^{\prime})\in D_{0}(f)$, consider the determinant $\Delta(\rho^{\prime})=\det M(\rho^{\prime})$ of the $3\times 3$ matrix $M(\rho^{\prime})$ which has as first row $x,y,z$, as second row $a,b,c$ and as third row $a^{\prime},b^{\prime},c^{\prime}$. Then it turns out that $\Delta(\rho^{\prime})$ is divisible by $f$, see [6]. For any multiple point $p\in{\mathcal{A}}$ we consider the corresponding local derivation $\tilde{D}_{p}$ and define a homogeneous polynomial $g_{p}\in S$ with $\deg g_{p}=d_{1}+1-m_{p}$ by setting
(1.4) $g_{p}=\frac{\Delta(\tilde{D}_{p})}{f}.$
These polynomials $g_{p}$ enter in the following description of free line arrangements.
Theorem 1.8.
With the above notation, let $m$ be the maximal multiplicity of a point in ${\mathcal{A}}$. Then the following hold.
  1. (1)
    If $m>d/2$, then $d_{1}=d-m$ and the line arrangement ${\mathcal{A}}$ is free if and only if
    $\tau({\mathcal{A}})=(d-1)^{2}-(d-m)(m-1).$
  2. (2)
    If $m\leq d/2$, then either $d_{1}=m-1$ and the arrangement ${\mathcal{A}}$ is free with exponents $d_{1}<d_{2}=d-m$, or
    $m\leq d_{1}\leq d-m.$
  3. (3)
    If $d-2\geq m\geq 2$, then the line arrangement ${\mathcal{A}}$ is free if and only if the ideal $I(\rho)\subset S$ spanned by all the polynomials $g_{p}$ with $p\in{\mathcal{A}}$ satisfying
    (1.5) $3\leq m_{p}\leq\frac{d+1}{2}$
    has an empty zero set in $\mathbb{P}^{2}$.
Note that the polynomial $g_{p}$ depends to some extent on the combinatorics, but not entirely, since we cannot describe $\rho$ in terms of the combinatorics, even when ${\mathcal{A}}$ is a free line arrangement. However, there is one situation where this can be done, and this leads to the following construction. Let ${\mathcal{A}}$ be a line arrangement in $\mathbb{P}^{2}$, and fix a point $q\in{\mathcal{A}}$, such that $m_{q}=m$ is the maximal multiplicity of points in ${\mathcal{A}}$. For any other point $p\neq q$, $p\in{\mathcal{A}}$ with $m_{p}\geq 3$, we define a line arrangement
${\mathcal{B}}_{p}:h_{p}=0$
as follows. If the points $p$ and $q$ are connected by a line $L\in{\mathcal{A}}$, then ${\mathcal{B}}_{p}$ consists simply of all the lines in ${\mathcal{A}}$ which contain neither $p$ nor $q$. If the points $p$ and $q$ are not connected by a line $L\in{\mathcal{A}}$, then ${\mathcal{B}}_{p}$ consists of all lines in ${\mathcal{A}}$ which contain neither $p$ nor $q$, plus the line $L(p,q)$ determined by the points $p$ and $q$. Note that in both cases one has
$|{\mathcal{B}}_{p}|=d-m_{p}-m_{q}+1.$
Proposition 1.9.
With the above notation, if the condition
$m>\frac{d}{2}$
is satisfied, then one may choose $\tilde{D}_{q}$ as the minimal degree derivation $\rho\in D_{0}(f)$ and with this choice one has
$g_{p}=h_{p}$
up to a nonzero constant factor, for any point $p$ distinct from $q$.
The polynomials $g_{p}$ introduced above enter also in the following result.
Theorem 1.10.
Let ${\mathcal{A}}:f=0$ be a line arrangement having the first exponent $d_{1}\geq 1$ and consider a point $p\in{\mathcal{A}}$ of multiplicity $m_{p}$ such that
$m_{p}\leq d_{1}+1<d-m_{p}+1.$
Then all the other multiple points of ${\mathcal{A}}$ which are not connected to $p$ by lines in ${\mathcal{A}}$ are situated on a curve ${\mathcal{C}}_{p}$ of degree at most $d_{1}+1-m_{p}$, which is a union of some irreducible components of the curve $g_{p}=0$. In particular, one has the following.
  1. (1)
    If $m_{p}=d_{1}+1$, then $p$ is a modular point and the line arrangement ${\mathcal{A}}$ is supersolvable.
  2. (2)
    If $m_{p}=d_{1}$, then all the other multiple points $q\in{\mathcal{A}}$ which are not connected to $p$ by lines in ${\mathcal{A}}$ are situated on a line $L$. If this line is not in ${\mathcal{A}}$, then
    $\sum_{q}m_{q}\leq d-m_{p},$
    where $m_{q}$ denote the multiplicity of point $q$ in ${\mathcal{A}}$. If this line is in ${\mathcal{A}}$, then one has the (usually) weaker inequality
    $\sum_{q}(m_{q}-1)\leq d-m_{p}-1.$
An alternative proof of the claim $(1)$ in Theorem 1.10 can be obtained using [7, Theorem 1.2] plus [9, Proposition 2.3]. For the claim $(2)$ we note that there are examples when the line $L$ cannot be chosen to be in ${\mathcal{A}}$, see Example 8.3.
Corollary 1.11.
Let ${\mathcal{A}}:f=0$ be a line arrangement having the first exponent $d_{1}\geq 1$. Then all the multiple points of ${\mathcal{A}}$ are situated on a curve ${\mathcal{C}}$ of degree at most $d_{1}+1$.
In fact, this result holds in a more general situation, when the line arrangement is replaced by an arrangement of smooth curves meeting transversally, see [24, Proposition 3.6] and [4, Remark 1.3]. Our proof shows that the curve ${\mathcal{C}}$ contains at least $m_{p}$ lines, so one may hope that the following more precise statement holds.
Conjecture 1.12.
Let ${\mathcal{A}}:f=0$ be a line arrangement in $\mathbb{P}^{2}$ having the first exponent $d_{1}\geq 1$. Let $N$ be the minimal number of lines in $\mathbb{P}^{2}$, such that all the multiple points of ${\mathcal{A}}$ are situated on a union of $N$ lines. Then
$N\leq d_{1}+1.$
As explained in the proof of Corollary 1.11, Conjecture 1.12 is equivalent to asking that the polynomial $g_{p}$ is a product of linear factors for at least one multiple point $p$ of any arrangement ${\mathcal{A}}$. Note that this happens for any $p\neq q$ in the conditions of Proposition 1.9 above, case which is discussed again in Theorem 1.14 $(3)$ below without the use of polynomials $g_{p}$. The next two results support Conjecture 1.12. The first one is a direct consequence of Theorem 1.10, $(2)$.
Corollary 1.13.
Let ${\mathcal{A}}:f=0$ be a line arrangement having the first exponent $d_{1}\geq 1$ and a point $p\in{\mathcal{A}}$ with multiplicity $m_{p}=d_{1}$. Then Conjecture 1.12 holds for the arrangement ${\mathcal{A}}$.
We have also the following.
Theorem 1.14.
Suppose that the line arrangement ${\mathcal{A}}$ has a point $p$ of multiplicity $m\geq 2$ such that $d_{1}=m-1$ or $d_{1}=d-m$, where $d=|{\mathcal{A}}|$. Then Conjecture 1.12 holds for the arrangement ${\mathcal{A}}$ and moreover the $N$ lines containing all the multiple points of ${\mathcal{A}}$ can be chosen in ${\mathcal{A}}$. In particular, this property holds for the following types of arrangements.
  1. (1)
    ${\mathcal{A}}$ has only double points and $d\geq 3$.
  2. (2)
    ${\mathcal{A}}$ has only double points except for $n_{3}$ triple points with $1\leq n_{3}\leq 3$ and $d\geq 4$.
  3. (3)
    ${\mathcal{A}}$ has a point $p$ of multiplicity $m$ such that $2m\geq|{\mathcal{A}}|$.
  4. (4)
    ${\mathcal{A}}$ is supersolvable.
Here is the organization of this paper. In Section 2 we recall some basic facts and then we prove Theorem 1.1 and Corollary 1.2. In Section 3 we give the proofs of Theorems 1.3 and 1.4. We also state and prove a related result involving $D_{0}(f)_{d-2}$, see Theorem 3.3 and Example 3.4.
In Section 4 we explain how the local derivations $\tilde{D}_{p}$ behave under addition-deletion mappings, and prove Theorem 1.6. In Section 5 we show first that the polynomial $g_{p}$ from (1.4) is nonzero when $\tilde{D}_{p}$ is not a constant multiple of minimal degree syzygy $\rho$, see Proposition 5.1. Then we prove Theorem 1.8 by using a key property of Bourbaki ideals associated with plane curves. Proposition 1.9 is also proved here.
In Section 6 we introduce the Euler type derivations, which lead to simpler formulas for the polynomials $g_{p}$ when the minimal degree derivation $D_{\rho}$ has this property. Using this approach, we give new characterizations for the monomial arrangement, see Theorem 6.5 and Corollary 6.12, and of the full monomial arrangement, see Theorem 6.8 and Corollary 6.9. Proposition 6.14 gives a necessary condition for a line arrangement to admit a minimal degree derivation of Euler type. In section 7 we prove some addition-deletion results for the Euler type derivations of line arrangements, see Propositions 7.1 and 7.3. In section 8 we prove Theorems 1.10 and 1.14 as well as Corollary 1.11
We would like to thank Takuro Abe, Dan Bath and Ştefan Tohăneanu for some useful discussions on the results in this paper.
2. The proof of Theorem 1.1 and of Corollary 1.2
We denote by $J_{f}$ the Jacobian ideal of $f$, i.e. the homogeneous ideal in $S$ spanned by $f_{x},f_{y},f_{z}$, and by $M(f)=S/J_{f}$ the corresponding graded ring, called the Jacobian (or Milnor) algebra of $f$. Let $I_{f}$ denote the saturation of the ideal $J_{f}$ with respect to the maximal ideal ${\bf m}=(x,y,z)$ in $S$ and consider the local cohomology group
$N(f)=I_{f}/J_{f}=H^{0}_{\bf m}(M(f)).$
The graded $S$-module $N(f)$ satisfies a Lefschetz type property with respect to multiplication by generic linear forms, see [12]. This implies in particular the inequalities
(2.1) $0\leq n(f)_{0}\leq...\leq n(f)_{\lceil T/2\rceil}\geq n(f)_{\lceil T/2\rceil+1}\geq...\geq n(f)_{T}\geq 0,$
where $T=3d-6$, $n(f)_{k}=\dim N(f)_{k}$ for any integer $k$, and for any real number $u$, $\lceil u\rceil$ denotes the round up of $u$, namely the smallest integer $U$ such that $U\geq u$. The self duality of the graded $S$-module $N(f)$, see [23], implies that
(2.2) $n(f)_{k}=n(f)_{T-k},$
for any integer $k$. The general form of the minimal resolution for the Milnor algebra $M(f)$ of such a curve $C:f=0$ is
(2.3) $0\to\displaystyle{\oplus_{i=1}^{m-2}S(-e_{i})}\to\displaystyle{\oplus_{i=1}^{m}S(1-d-d_{i})}\to S^{3}(1-d)\to S,$
with $e_{1}\leq e_{2}\leq\ldots\leq e_{m-2}$ and $1\leq d_{1}\leq d_{2}\leq\cdots\leq d_{m}$. It follows from [19, Lemma 1.1] that one has
$e_{j}=d+d_{j+2}-1+\epsilon_{j},$
for $j=1,\ldots,m-2$ and some integers $\epsilon_{j}\geq 1.$ The minimal resolution of $N(f)$ obtained from (2.3), by [19, Proposition 1.3], is
$0\to\displaystyle{\oplus_{i=1}^{m-2}S(-e_{i})}\to\displaystyle{\oplus_{i=1}^{m}S(-\ell_{i})}\to\displaystyle{\oplus_{i=1}^{m}}S(d_{i}-2(d-1))\to\displaystyle{\oplus_{i=1}^{m-2}S(e_{i}-3(d-1))},$
where $\ell_{i}=d+d_{i}-1$. It follows that
(2.4) $\sigma(C)=3(d-1)-e_{m-2}=2(d-1)-d_{m}-\epsilon_{m-2},$
where
$\sigma(C)=\min\left\{j:n(f)_{j}\neq 0\right\}=\operatorname{indeg}(N(f)).$
2.1. The proof of Theorem 1.1
Using [6, Equation (2.3)], we get that for any integer $k$ one has
$\dim D_{0}(f)_{k+1}-\dim N(f)_{d+k}+\dim D_{0}(f)_{d-5-k}=$
$=\tau({\mathcal{A}})+3\binom{k+3}{2}-\binom{d+k+2}{2}.$
If $k\geq d-4$, then $D_{0}(f)_{d-5-k}=0$ since $d-5-k<0$. Next we show that
(2.5) $N(f)_{2d-4}=0,$
or, equivalently, $N(f)_{d-2}=0$, recall (2.2). In view of (2.2), this is equivalent to showing
$\sigma({\mathcal{A}})\geq d-1.$
Note that the minimal resolution (2.3) gives rise to the following minimal resolution for $D_{0}(f)$:
(2.6) $0\to\displaystyle{\oplus_{j=1}^{m-2}S(-(d_{j+2}+\epsilon_{j}))}\to\displaystyle{\oplus_{i=1}^{m}S(-d_{i})}.$
It is known that the Castelnuovo-Mumford regularity of the $S$-module $D_{0}({\mathcal{A}})$ is at most $d-2$, see [22, Corollary 3.5] and also [11]. This implies that
(2.7) $d_{m}\leq d-2\text{ and }d_{m}+\epsilon_{m-2}-1\leq d-2.$
Using the second inequality we get
$\sigma({\mathcal{A}})=2(d-1)-d_{m}-\epsilon_{m-2}\geq d-1,$
which is exactly what we need to get (2.5). Now, for $k\geq d-4$, we have $d+k\geq 2d-4$ and hence $N(f)_{d+k}=0$. Indeed, we have shown that $N(f)_{2d-4}=0$ and the claim follows using (2.1) since
$T/2=3(d-2)/2\leq 2d-4$
for $d\geq 2$. This completes the proof of claim $(1)$.
To prove $(2)$, recall the Ziegler pair of two arrangements ${\mathcal{A}}:f=0$ and ${\mathcal{A}}^{\prime}:f^{\prime}=0$ of $d=9$ lines, such that $D_{0}(f)_{5}\neq 0$ and $D_{0}(f^{\prime})_{5}=0$, see [25, 16].
2.2. The proof of Corollary 1.2
Use Theorem 1.1 and the following two well known formulas
$\tau({\mathcal{A}})=\sum_{p}(m_{p}-1)^{2}\text{ and }\sum_{p}\binom{m_{p}}{2}=\binom{d}{2}.$
3. The proofs of Theorems 1.3 and 1.4 and a related result
3.1. The proof of Theorem 1.3
To prove $(1)$, it is enough that for any line $L:\ell=0$ in $\mathbb{P}^{2}$, one has that $p\in L$ if and only if $D_{p}(\ell)=0$. Indeed, $f_{1p}$ is just a product of such linear forms $\ell$ with $D_{p}(\ell)=0$.
One can prove $(2)$ either by a direct computation or using [13, Theorem 5.1].
To prove $(3)$, we compute the coefficients $A,B,C$ of ${\partial_{x}}$, ${\partial_{y}}$ and respectively ${\partial_{z}}$ in $\tilde{D}_{p}(p)$ and we get
$(A:B:C)=\frac{m_{p}}{d}f_{2p}(p)(u:v:w),$
since $D_{p}(f_{2p})(p)=d_{2}f_{2p}(p)$, with $d_{2}=\deg f_{2p}=d-m_{p}>0$.
To prove $(4)$, note that $f_{2p}$ vanishes at $q$ of order $m_{q}$ (resp. $m_{q}-1$) if the line $L(p,q)$ determined by $p$ and $q$ is not in ${\mathcal{A}}$ (resp. if $L(p,q)$ is in ${\mathcal{A}}$). Hence $f_{2p}$ and $D_{p}(f_{2p})$ vanish at $q$ of order at least $m_{q}-2$.
Finally, to prove $(5)$, note that we may suppose $p=(1:0:0)$. With this choice one has
$\tilde{D}_{p}=\left(f_{2p}-\frac{xf_{2p,x}}{d}\right)\partial_{x}-\frac{yf_{2p,x}}{d}\partial_{y}-\frac{zf_{2p,x}}{d}\partial_{z}.$
A point $q=(u:v:w)\neq p$ has either $v\neq 0$ or $w\neq 0$. It follows that $\tilde{D}_{p}(q)=0$ if and only if $f_{2p,x}(q)=f_{2p}(q)=0$, in particular $q\in{\mathcal{A}}_{2}$. If $q$ is a point of multiplicity at least two in ${\mathcal{A}}_{2}$, then clearly $\tilde{D}_{p}(q)=0$, since $f_{2p}$ and all of its first derivatives vanish at $q$. If $q$ is a simple point of ${\mathcal{A}}_{2}$, and $L:\ell=0$ is the unique line in ${\mathcal{A}}_{2}$ containing $q$, then one has
$(f_{2p,x}(q):f_{2p,y}(q):f_{2p,z}(q))=({\alpha}:{\beta}:\gamma),$
where $\ell={\alpha}x+{\beta}y+\gamma z$. Indeed, both sides give the coefficients of the tangent line $L:\ell=0$ to ${\mathcal{A}}_{2}$ at the simple point $q$. It follows that ${\alpha}=0$ and hence $L$ is a line containing $p$, a contradiction, since $L$ was supposed to be a line in ${\mathcal{A}}_{2}$. Hence this second case is impossible, and this proves the claim $(5)$.
3.2. The proof of Theorem 1.4
We show that there is no notrivial relation
$\sum_{p}P_{p}\tilde{D}_{p}=0,$
where $P_{p}\in S_{m_{p}-3}$ are not all zero. Let $M$ be a minimal subset of the set of points in ${\mathcal{A}}$ of multiplicity at least $3$ such that
$\sum_{p\in M}P_{p}\tilde{D}_{p}=0,$
where $P_{p}\in S_{m_{p}-3}$ are nozero for any $p\in M$. Let $q\in M$ be a point of minimal multiplicity, that is $m_{q}\leq m_{p}$ for all $p\in M$. Then using Theorem 1.3, claims $(3)$ and $(4)$, we see that the coefficient $P_{q}$ has to vanish at $q$ of order at least $m_{q}-2$. Indeed, all the other coefficients of $\tilde{D}_{p}$ for $p\neq q$ have this property, and at least one coefficient of $\tilde{D}_{q}$ does not vanish at $q$. Since $P_{q}$ is a homogeneous polynomial of degree $m_{q}-3$, this implies $P_{q}=0$ a contradiction.
We have also the following result.
Theorem 3.3.
Let $p,p^{\prime}$ and $p^{\prime\prime}$ be three non-collinear multiple points in an arrangement ${\mathcal{A}}$ of $d$ lines, such that at most two of the three lines determined by $p$, $p^{\prime}$ and $p^{\prime\prime}$ are in ${\mathcal{A}}$. Then $D_{0}(f)_{d-2}$ contains the direct sum
$S_{m-2}\tilde{D}_{p}\oplus S_{m^{\prime}-2}\tilde{D}_{p^{\prime}}\oplus S_{m^{\prime\prime}-2}\tilde{D}_{p^{\prime\prime}},$
where $m=\operatorname{mult}_{p}({\mathcal{A}})\geq 3$, $m^{\prime}=\operatorname{mult}_{p^{\prime}}({\mathcal{A}})\geq 3$ and $m^{\prime\prime}=\operatorname{mult}_{p^{\prime\prime}}({\mathcal{A}})\geq 3$.
Proof.
Since $p,p^{\prime}$ and $p^{\prime\prime}$ are non-collinear, we may choose the coordinates such that
$p=(1:0:0),\ p^{\prime}=(0:1:0)\text{ and }p^{\prime\prime}=(0:0:1).$
With this choice, we get 3 derivatives as in Theorem 1.3, namely
$\tilde{D}_{p}=f_{2p}\partial_{x}-\frac{f_{2p,x}}{d}E,\ \tilde{D}_{p^{\prime}}=f_{2p^{\prime}}\partial_{y}-\frac{f_{2p^{\prime},y}}{d}E\text{ and }\tilde{D}_{p^{\prime\prime}}=f_{2p^{\prime\prime}}\partial_{z}-\frac{f_{2p^{\prime\prime},z}}{d}E.$
Let $P,P^{\prime}$ and $P^{\prime\prime}$ be homogeneous polynomials in $S$ of degree $m-2,\ m^{\prime}-2$ and respectively $m^{\prime\prime}-2$. We show that
$\Sigma=P\tilde{D}_{p}+P^{\prime}\tilde{D}_{p^{\prime}}+P^{\prime\prime}\tilde{D}_{p^{\prime\prime}}=0$
implies $P=P^{\prime}=P^{\prime\prime}=0$. By symmetry, we can assume that the line $x=0$ is not in ${\mathcal{A}}$. Let $A$ be the coefficient of $\partial_{x}$ in the sum $\Sigma\in D_{0}(f)_{d-2}$. We get
$dA=P(df_{2p}-xf_{2p,x})-P^{\prime}xf_{2p^{\prime},y}-P^{\prime\prime}xf_{2p^{\prime\prime},z}.$
Since $x=0$ is not in ${\mathcal{A}}$, it follows that $f_{2p}$ is not divisible by $x$. Then $A=0$ yields $P=xP_{1}$, for some homogeneous polynomial $P_{1}$ of degree $m-3$. Note that $\deg f_{2p}=d-m$ and $f_{2p}(p)\neq 0$ implies that $x^{d-m}$ occurs in $f_{2p}$ with a non-zero coefficient. It follows that the factor $(df_{2p}-xf_{2p,x})$ does not vanish at $p$. Now note that $f_{2p^{\prime}}$ and $f_{2p^{\prime\prime}}$ both vanish at $p$ of order at least $m-1$. More precisely, $f_{2p^{\prime}}$ vanishes at $p$ of order $m-1$ if the line $z=0$ is in ${\mathcal{A}}$, and of order $m$ if the line $z=0$ is not in ${\mathcal{A}}$. Hence $A=0$ would imply that $P_{1}$ vanishes at $p$ of order at least $m-2$. Since $\deg P=m-3$, this is possible only if $P_{1}=0$. Hence $P=0$, and we are left with a relation
$\Sigma_{1}=P^{\prime}\tilde{D}_{p^{\prime}}+P^{\prime\prime}\tilde{D}_{p^{\prime\prime}}=0.$
Moreover $A=0$ and $P=0$ give
$P^{\prime}f_{2p^{\prime},y}+P^{\prime\prime}f_{2p^{\prime\prime},z}=0$
Let $B$ be the coefficient of $\partial_{y}$ in the sum $\Sigma_{1}$. Then we get
$dB=P^{\prime}(df_{2p^{\prime}}-yf_{2p^{\prime},y})-P^{\prime\prime}yf_{2p^{\prime\prime},z}=0=dP^{\prime}f_{2p^{\prime}}-y(P^{\prime}f_{2p^{\prime},y}+P^{\prime\prime}f_{2p^{\prime\prime},z})=dP^{\prime}f_{2p^{\prime}}=0.$
This implies $P^{\prime}=0$ and then of course $P^{\prime\prime}=0$ as well. This ends the proof of our claim.
Example 3.4.
The vector space $D_{0}(f)_{d-2}$ does not have a the direct sum decomposition similar to that in Theorem 1.4 and the extra assumption in Theorem 3.3 is necessary. Consider for instance the line arrangement
${\mathcal{A}}:xyz(x-y)(y-z)(x-z)=0.$
We consider the triple points of ${\mathcal{A}}$, that is the points
$q=(1:1:1),\ p=(1:0:0),\ p^{\prime}=(0:1:0)\text{ and }p^{\prime\prime}=(0:0:1).$
It is known that ${\mathcal{A}}$ is free with exponents $(2,3)$ and hence
$\dim D_{0}(f)_{4}=\dim S_{2}+\dim S_{1}=6+3=9.$
On the other hand
$S_{1}\tilde{D}_{p}+S_{1}\tilde{D}_{p^{\prime}}+S_{1}\tilde{D}_{p^{\prime\prime}}\neq D_{0}(f)_{4},$
since $\tilde{D}_{p}(q)=\tilde{D}_{p^{\prime}}(q)=\tilde{D}_{p^{\prime\prime}}(q)=0$ by Theorem 1.3 $(4)$ and $D=x\tilde{D}_{q}\in D_{0}(f)_{4}$ satisfies $D(q)\neq 0$ by Theorem 1.3 $(3)$. It follows that there are non-trivial relations $\Sigma=0$ as in the proof of Theorem 1.4, with $\deg P=\deg P^{\prime}=\deg P^{\prime\prime}=1$. In other words, the condition that at most two of the three lines determined by $p$, $p^{\prime}$ and $p^{\prime\prime}$ are in ${\mathcal{A}}$ is necessary for our proof of Theorem 3.3 to hold.
4. Addition-deletion for local derivations and the proof of Theorem 1.6
4.1. Addition-deletion for local derivations
In this section we assume that ${\mathcal{A}}={\mathcal{A}}^{\prime}\cup L$, where ${\mathcal{A}}^{\prime}:f^{\prime}=0$ is line arrangement obtained by deleting the line $L$ from ${\mathcal{A}}$. Let ${\mathcal{A}}^{\prime\prime}:f^{\prime\prime}=0$ be the multiarrangement on $L$ defined by the restriction of $f$ to $L$. We choose the coordinates such that $L:x=0$ and, then $f=xf^{\prime}$ and $f^{\prime\prime}=\overline{f}^{\prime}$, where we define for any polynomial $g\in S$
$\overline{g}(y,z)=g(0,y,z).$
In this setting there is an exact sequence of graded $S$-modules
(4.1) $0\to D(f^{\prime})(-1)\stackrel{{\scriptstyle u}}{{\longrightarrow}}D(f)\stackrel{{\scriptstyle v}}{{\longrightarrow}}D(f^{\prime\prime}),$
where the morphisms $u$ and $v$ are given by
(4.2) $u(\delta)=x\delta\text{ and }v(a,b,c)=(\overline{b},\overline{c})$
and $D(f^{\prime\prime})$ is regarded as an $S$-module via the morphism $S\to\mathbb{C}[y,z]$, $g\mapsto\overline{g}$, see [21, Proposition 4.45]. One may ask whether a similar exact sequence exists for the modules $D_{0}(f^{\prime})$, $D_{0}(f)$ and $D_{0}(f^{\prime\prime})$, namely
(4.3) $0\to D_{0}(f^{\prime})(-1)\stackrel{{\scriptstyle u_{0}}}{{\longrightarrow}}D_{0}(f)\stackrel{{\scriptstyle v_{0}}}{{\longrightarrow}}D_{0}(f^{\prime\prime}).$
This answer is known to be positive and can be recovered from [22], where $D_{0}(f^{\prime})$, $D_{0}(f)$ and $D_{0}(f^{\prime\prime})$ are identified to $D(f^{\prime})/S\cdot E$, $D(f)/S\cdot E$ and respectively to $D(f^{\prime\prime})/S\cdot\overline{E}$. This identification gives simpler, but less precise formulas for $u_{0}$ and $v_{0}$. It is easy to see by a direct computation that the following result holds. Details for the proof in a more general setting will be given elsewhere.
Theorem 4.2.
With the above notation, the sequence in (4.3) is exact if we define the morphisms $u_{0}$ and $v_{0}$ of graded $S$-modules as follows.
  1. (1)
    For $\delta=(a,b,c)\in D_{0}(f^{\prime})$, we set
    $u_{0}(\delta)=x\delta-\frac{aE}{d}$
    with $d=\deg f$ and $E=(x,y,z)$ the Euler derivation,
  2. (2)
    Any $\delta\in D_{0}(f)$ has the form $\delta=(xa_{1},b,c)$ with $a_{1}\in S$, and with this notation we set
    $v_{0}(\delta)=(\overline{b},\overline{c})+\frac{\overline{a}_{1}}{d-1}(y,z).$
In order to use Theorem 4.2, we first see how the local derivations $\tilde{D}_{p}$ are mapped under the morphisms $u_{0}$ and $v_{0}$.
Proposition 4.3.
Consider an arrangement ${\mathcal{A}}$ of $d$ lines in $\mathbb{P}^{2}$ such that $L:x=0$ is a line in ${\mathcal{A}}$ and set ${\mathcal{A}}^{\prime}={\mathcal{A}}\setminus\{L\}$. For a multiple point $p\in{\mathcal{A}}^{\prime}$ we denote by $\tilde{D}_{p}^{\prime}\in D_{0}({\mathcal{A}}^{\prime})$ the local derivation associated to $p$, and by $\tilde{D}_{p}\in D_{0}({\mathcal{A}})$ the local derivation associated to $p\in{\mathcal{A}}$. Then the following two cases are possible.
  1. (1)
    If $p\in L$, then $u_{0}(\tilde{D}_{p}^{\prime})=x\tilde{D}_{p}$.
  2. (2)
    If $p\notin L$, then $u_{0}(\tilde{D}_{p}^{\prime})=\tilde{D}_{p}$.
Proof.
To prove the claim $(1)$, we may assume that $p=(0:1:0)$ and $f=xf^{\prime}$. Then, with obvious notation, we have $f_{1}=xf_{1}^{\prime}$ and $f_{2}=f_{2}^{\prime}$. It follows that
$\tilde{D}_{p}^{\prime}=f_{2}\partial_{y}-\frac{f_{2,y}}{d-1}E$
and hence
$u_{0}(\tilde{D}_{p}^{\prime})=x\tilde{D}_{p}^{\prime}+\frac{xf_{2,y}}{d(d-1)}E=x\tilde{D}_{p}.$
To prove the claim $(2)$, we may assume that $p=(1:0:0)$ and $f=xf^{\prime}$. In this case we have $f_{1}=f_{1}^{\prime}$ and $f_{2}=xf_{2}^{\prime}$. It follows that
$\tilde{D}_{p}^{\prime}=f_{2}^{\prime}\partial_{x}-\frac{f_{2,x}^{\prime}}{d-1}E$
and hence
$u_{0}(\tilde{D}_{p}^{\prime})=x\tilde{D}_{p}^{\prime}-\frac{(d-1)f_{2}^{\prime}-xf_{2,x}}{d(d-1)}E=\tilde{D}_{p}.$
Proposition 4.4.
Consider an arrangement ${\mathcal{A}}$ of $d$ lines in $\mathbb{P}^{2}$ such that $L:x=0$ is a line in ${\mathcal{A}}$ and set ${\mathcal{A}}^{\prime\prime}:f^{\prime\prime}=0$ to be the multiarrangement obtained by restriction of ${\mathcal{A}}$ to $L$. For a multiple point $p\in{\mathcal{A}}$ we denote by $\tilde{D}_{p}\in D_{0}({\mathcal{A}})$ the local derivation associated to $p$. Then the following two cases are possible.
  1. (1)
    If $p=(0:-\gamma_{p}:{\beta}_{p})\in L$, then we set $f=f_{1}f_{2}$ and $f^{\prime\prime}=\overline{f}_{2}=f_{2}(0,y,z)$ as above. With this notation, one has
    $v_{0}(\tilde{D}_{p})=\frac{1}{d-1}(yA_{p}-(d-1)\gamma_{p}\overline{f}_{2},zA_{p}+(d-1){\beta}_{p}\overline{f}_{2}),$
    where $A_{p}=\gamma_{p}\overline{f}_{2,y}-{\beta}_{p}\overline{f}_{2,z}$.
  2. (2)
    If $p\notin L$, then $v_{0}(\tilde{D}_{p})=0$.
Proof.
To prove the first claim, notice that
$\tilde{D}_{p}=f_{2}(0,-\gamma_{p},{\beta}_{p})+\frac{A_{p}}{d}(x,y,z).$
The formula for $v_{0}(\tilde{D}_{p})$ follows now from the definition of $v_{0}$ in Theorem 4.2. The second claim follows from the exact sequence in Theorem 4.2 and Proposition 4.3, claim $(2)$. ∎
4.5. The proof of Theorem 1.6
With the notation from Propositions 4.4, let $K=L\cap{\mathcal{A}}^{\prime}$ and let $k=|K|$ be the number of points in this intersection. Since ${\mathcal{A}}$ is not a pencil, we get $k\geq 2$. Let $p_{1},\ldots,p_{k}$ be the points in $K$ and assume that
$p_{i}=(0:-\gamma_{i}:{\beta}_{i})\text{ for }i=1,\ldots,k.$
Then one has $f=xf^{\prime}$ and
(4.4) $f^{\prime\prime}(y,z)=f^{\prime}(0,y,z)=\prod_{i=1}^{k}\ell_{i}^{m_{i}-1},$
where $\ell_{i}={\beta}_{i}y+\gamma_{i}z$ and $m_{i}=m_{p_{i}}$. In particular, we have
(4.5) $\sum_{i=1}^{k}(m_{i}-1)=d-1.$
It follows that the $S$-module $D_{0}(f^{\prime\prime})$ has one generator $\delta^{\prime\prime}$, namely the one obtained from the Koszul derivation
$\delta_{K}=f^{\prime\prime}_{z}\partial_{y}-f"_{y}\partial_{z}$
by removing the common factor
$G.C.D.(f^{\prime\prime}_{y},f^{\prime\prime}_{z})=\prod_{i=1}^{k}\ell_{i}^{m_{i}-2}.$
It follows that
(4.6) $\delta^{\prime\prime}=\sum_{i=1}^{k}(m_{i}-1)(\prod_{s\neq i}\ell_{s})(\gamma_{i}\partial_{y}-{\beta}_{i}\partial_{z})=B^{\prime\prime}\partial_{y}+C^{\prime\prime}\partial_{z}.$
Now we consider the derivation $\tilde{D}_{p_{1}}$ and find out the relation between $v_{0}(\tilde{D}_{p_{1}})$ and $\delta^{\prime\prime}$. To use Proposition 4.4 $(1)$, we first notice that all the lines in ${\mathcal{A}}$ which do not pass through $p_{1}$, intersect $L$ in the points $p_{s}$ with $s\neq 1$. It follows that, if $f_{2}=0$ is the equation of the union of these lines in ${\mathcal{A}}$, one has
(4.7) $g_{1}=f_{2}(0,y,z)=\prod_{s=1,\ldots,k;s\neq 1}\ell_{s}^{m_{s}-2},$
where the index $1$ in $g_{1}$ says that this polynomial is constructed with respect to the point $p_{1}\in K$. Proposition 4.4 $(1)$ yields the following equality, where the constant factor $1/(d-1)$ was omitted since it plays no role
$v_{0}(\tilde{D}_{p_{1}})=g_{1}(B_{1}\partial_{y}+C_{1}\partial_{z})$
where
$B_{1}=\gamma_{1}y\sum_{s\neq 1}\left((m_{s}-1){\beta}_{s}\prod_{t\neq 1,t\neq s}\ell_{t}\right)-{\beta}_{1}y\sum_{s\neq 1}\left((m_{s}-1)\gamma_{s}\prod_{t\neq 1,t\neq s}\ell_{t}\right)-(d-1)\gamma_{1}\prod_{s\neq 1}\ell_{s}$
and
$C_{1}=\gamma_{1}z\sum_{s\neq 1}\left((m_{s}-1){\beta}_{s}\prod_{t\neq 1,t\neq s}\ell_{t}\right)-{\beta}_{1}z\sum_{s\neq 1}\left((m_{s}-1)\gamma_{s}\prod_{t\neq 1,t\neq s}\ell_{t}\right)+(d-1){\beta}_{1}\prod_{s\neq 1}\ell_{s},$
where $s,t$ take all values from $1$ to $k$ except those indicated in the sums and respectively in the products above. Using (4.5) we can rewrite $B_{1}$ as follows
$B_{1}=\gamma_{1}y\sum_{s\neq 1}\left((m_{s}-1){\beta}_{s}\prod_{t\neq 1,t\neq s}\ell_{t}\right)-{\beta}_{1}y\sum_{s\neq 1}\left((m_{s}-1)\gamma_{s}\prod_{t\neq 1,t\neq s}\ell_{t}\right)-$
$-\gamma_{1}\sum_{s\neq 1}\left((m_{s}-1)\ell_{s}\prod_{t\neq 1,t\neq s}\ell_{t}\right)-\gamma_{1}(m_{1}-1)\prod_{s\neq 1}\ell_{s}.$
Note that
$\gamma_{1}{\beta}_{s}y-\gamma_{1}\ell_{s}=-\gamma_{1}\gamma_{s}z.$
Hence after cancelation of terms in the first and third sum above we get
$B_{1}=-\ell_{1}\sum_{s\neq 1}\left((m_{s}-1)\gamma_{s}\prod_{t\neq 1,t\neq s}\ell_{t}\right)-\gamma_{1}(m_{1}-1)\prod_{s\neq 1}\ell_{s}.$
It follows that $B_{1}=-B^{\prime\prime}$, and a similar computation yields $C_{1}=-C^{\prime\prime}$. In conclusion, one has
(4.8) $v_{0}(\tilde{D}_{p_{1}})=g_{1}(B_{1}\partial_{y}+C_{1}\partial_{z})=-g_{1}\delta^{\prime\prime}.$
By symmetry, if we set for any $i\in[2,k]$,
(4.9) $g_{i}=\prod_{s=1,\ldots,k;s\neq i}\ell_{s}^{m_{s}-2},$
we get the equality
(4.10) $v_{0}(\tilde{D}_{p_{i}})=-g_{i}\delta^{\prime\prime},$
where as above the constant factor $1/(d-1)$ is omitted. To continue the proof of Theorem 1.6, we show an inclusion of ideals in the ring $\mathbb{C}[y,z]$, namely
(4.11) $(y,z)^{d-2-k}\subset I=(g_{1},\ldots,g_{k}).$
Note that in view of (4.5) we have.
$\sum_{i=1}^{k}(m_{i}-2)=d-1-k$
and that we may assume $\ell_{1}=y$ and $\ell_{2}=z$. We prove this claim by induction on $k$. If $k=2$, the claim is clear. Indeed, any monomial $y^{s}z^{t}$ of degree $s+t=d-4$ cannot satisfy
$s\leq m_{1}-3\text{ and }t\leq m_{2}-3,$
since then $s+t\leq(m_{1}-3)+(m_{2}-3)=d-5$, a contradiction. Hence $y^{s}z^{t}$ is divisible either by $g_{2}$ or by $g_{1}$, and hence $y^{s}z^{t}\in I$ in this case. Assume now the claim (4.11) proved for any set of $k-1$ linear forms. By this hypothesis, it follows that we have the inclusions of ideals
$y^{m_{1}-2}(y,z)^{\sum_{i\neq 1}(m_{i}-2)-1}\subset I\text{ and }z^{m_{2}-2}(y,z)^{\sum_{i\neq 2}(m_{i}-2)-1}\subset I.$
If we have a monomial $y^{s}z^{t}$ such that $s+t=d-2-k$, then two cases are possible. Either $s\geq m_{1}-2$ and then
$y^{s}z^{t}\in y^{m_{1}-2}(y,z)^{\sum_{i\neq 1}(m_{i}-2)-1},$
or $t\geq m_{2}-2$ and then
$y^{s}z^{t}\in z^{m_{2}-2}(y,z)^{\sum_{i\neq 2}(m_{i}-2)-1}.$
Indeed, if $s\leq m_{1}-3$ and $t\leq m_{2}-3$ we get
$\sum_{i=1}^{k}(m_{i}-2)=d-1-k=s+t+1\leq(m_{1}-2)+(m_{2}-2)-1.$
This is a contradiction, since $m_{i}-2\geq 0$ for any $i$. Hence the proof of the inclusion (4.11) is completed.
Now we prove Theorem 1.6 by induction on $d=|{\mathcal{A}}|\geq 3$. Since we assume that ${\mathcal{A}}$ is not a pencil, the starting point of the induction is ${\mathcal{A}}:f=xyz=0$. Then it is known that ${\mathcal{A}}$ is free with exponents $(1,1)$. In particular $\dim D_{0}({\mathcal{A}})_{1}=2$ and any basis $\rho_{1},\rho_{2}$ of the vector space $D_{0}({\mathcal{A}})_{1}$ freely generate the $S$-module $D_{0}({\mathcal{A}})$. For the double point $p=(1:0:0)$ (resp. $p^{\prime}=(0:1:0)$) we get
$\tilde{D}_{p}=x\partial_{x}-\frac{1}{3}E\in D_{0}({\mathcal{A}})_{1}$
and respectively
$\tilde{D}_{p^{\prime}}=y\partial_{y}-\frac{1}{3}E\in D_{0}({\mathcal{A}})_{1}.$
These two local derivations are clearly independent, form a basis of the vector space $D_{0}({\mathcal{A}})_{1}$, and hence generate $D_{0}({\mathcal{A}})_{k}$ for any $k\geq d-2=1$. Assume the claim in Theorem 1.6 holds for all arrangements ${\mathcal{A}}^{\prime}$ with $|{\mathcal{A}}^{\prime}|<d$ and consider the exact sequence (4.3), which yields
(4.12) $0\to D_{0}(f^{\prime})_{d-3}\stackrel{{\scriptstyle u_{0}}}{{\longrightarrow}}D_{0}(f)_{d-2}\stackrel{{\scriptstyle v_{0}}}{{\longrightarrow}}D_{0}(f^{\prime\prime})_{d-2}.$
By the induction hypothesis, $D_{0}(f^{\prime})_{d-3}$ is generated by local derivations corresponding to the multiple points of ${\mathcal{A}}^{\prime}$. Proposition 4.3 says that $\operatorname{im}u_{0}$ is contained in the graded submodule $M\subset D_{0}({\mathcal{A}})$ generated by the local derivations corresponding to the multiple points of ${\mathcal{A}}$. On the other hand, the equalities (4.10) and the inclusion (4.11) imply that the morphism
$M_{d-2}\stackrel{{\scriptstyle v_{0}}}{{\longrightarrow}}D_{0}(f^{\prime\prime})_{d-2}=S_{d-k-1}\rho^{\prime\prime}$
obtained by restriction of $v_{0}$ to $M_{d-2}$ is surjective. The exact sequence (4.12) shows that $M_{d-2}=D_{0}(f)_{d-2}$, which is enough to prove Theorem 1.6 in view of our Remark 1.7 above.
5. Local derivations, Bourbaki ideals and the proof of Theorem 1.8 and Proposition 1.9
We recall first the construction of the Bourbaki ideal $B(C,\rho_{1})$ associated to a degree $d$ reduced curve $C:f=0$ and to a minimal degree non-zero syzygy $\rho\in D_{0}(f)$, see [15] as well as [20] for a recent, more complete approach. For any choice of the syzygy $\rho=(a,b,c)$ with minimal degree $d_{1}$, we have a morphism of graded $S$-modules
[…]
For any homogeneous syzygy $\rho^{\prime}=(a^{\prime},b^{\prime},c^{\prime})\in D_{0}(f)_{m}$, consider the determinant $\Delta(\rho^{\prime})=\det M(\rho^{\prime})$ of the $3\times 3$ matrix $M(\rho^{\prime})$ which has as first row $x,y,z$, as second row $a,b,c$ and as third row $a^{\prime},b^{\prime},c^{\prime}$. Then it turns out that $\Delta(\rho^{\prime})$ is divisible by $f$, see [6], and we define thus a new morphism of graded $S$-modules
[…]
and a homogeneous ideal $B(C,\rho)\subset S$ such that $\operatorname{im}v=B(C,\rho)(d_{1}-d+1)$. We have
(5.3) $B(C,\rho)=S\text{ if and only if }C\text{ is a free curve}.$
It is known that the ideal $B(C,\rho)$, when $C$ is not a free curve, defines a 00-dimensional subscheme in $\mathbb{P}^{2}$, which is locally a complete intersection, see for details [15, Theorem 5.1].
Let ${\mathcal{A}}:f=0$ be a line arrangement and fix a minimal degree non-zero syzygy $\rho\in D_{0}(f)$, say $\rho=(a,b,c)$ and $\deg\rho=d_{1}$. For any multiple point $p\in{\mathcal{A}}$ we consider the corresponding local derivation $\tilde{D}_{p}$ and define a polynomial $g_{p}\in S$, such that one has
$g_{p}=v(\tilde{D}_{p})\in B({\mathcal{A}},\rho).$
Using the formula for $\tilde{D}_{p}$ given in Theorem 1.3 we see that
$\Delta(\tilde{D}_{p})=f_{2p}\Delta_{p}$
where $\Delta_{p}$ is the determinant of the $3\times 3$ matrix $M_{p}$ which has as first row $x,y,z$, as second row $a,b,c$ and as third row $u,v,w$ if $p=(u:v:w)$. It follows that
(5.4) $g_{p}=\frac{\Delta(\tilde{D}_{p})}{f}=\frac{\Delta_{p}}{f_{1p}}.$
One has the following result.
Proposition 5.1.
With the above notation, for each multiple point $p\in{\mathcal{A}}$ of multiplicity $m_{p}$ satisfying (1.5) the corresponding polynomial $g_{p}\in S$ is homogeneous of degree $d_{1}+1-m_{p}\geq 0$ when $\tilde{D}_{p}$ is not a constant multiple of $\rho$. In particular $g_{p}\neq 0$ in this situation.
Proof.
Note that $g_{p}=0$ is equivalent to $\Delta(\tilde{D}_{p})=0$. This last condition is equivalent to the fact that the rows in the matrix $M(\Delta(\tilde{D}_{p}))$ are linearly dependent over the field of franctions of the polynomial ring $S$. This occurs if and only if there is a nontrivial relation
$PE+QD_{\rho}+R\tilde{D}_{p}=0,$
where $P,Q,R\in S$ and $E$ is as above the Euler derivation. If we apply this relation to $f$, we get $dPf=0$, and hence $P=0$. Suppose $\tilde{D}_{p}$ is not a constant multiple of the minimal degree derivation $\rho$. Since $\tilde{D}_{p}$ is a primitive derivation, it follows that $\tilde{D}_{p}$ is not a polynomial multiple of $\rho$. Then the result follows, since the quotient $D_{0}(f)/\operatorname{im}u$, with $u$ as in (5.1), is a torsion free $S$-module, being isomorphic to $\operatorname{im}v$. Hence $Q=R=0$, a contradiction which implies that $g_{p}\neq 0$. If $g_{p}\neq 0$, then the claim $\deg g_{p}=d_{1}+1-m_{p}$ follows from (5.4). ∎
5.2. Proof of Theorem 1.8
Let $q\in{\mathcal{A}}$ be a point of multiplicity $m$. If $m>d/2$, then the associated derivation $\tilde{D}_{q}$ is a derivation in $D_{0}(f)$ of minimal degree $r=d-m\leq(d-1)/2$, see [7, Theorem 1.2]. The claim about freeness of ${\mathcal{A}}$ follows from [17, 6]. This proves the claim $(1)$. To prove the claim $(2)$, we apply again [7, Theorem 1.2] and obtain the claim for the case $d_{1}=m-1$, as well as the inequalities $m\leq d_{1}\leq d-m$.
The proof for $(3)$ is more subtle. The case $m=2$ is clear, since we have $d\geq 4$ and such a nodal arrangement is never free. On the other hand $I(\rho)=0$ in this case. Assume from now on that $m\geq 3$. Consider first the case when ${\mathcal{A}}$ is free, with exponents $d_{1}\leq d_{2}$. Then $d_{1}\geq 2$ by $(1)$ and $(2)$ and hence
$d_{2}=d-1-d_{1}\leq d-3.$
It follows that the second generator $\rho_{2}$ of $D_{0}(f)$ has degree $d_{2}\leq d-3$ and hence we have the equality
$S_{d-3-d_{1}}\rho_{1}+S_{d-3-d_{2}}\rho_{2}=D_{0}(f)_{d-3}.$
On the other hand, since ${\mathcal{A}}$ is free, we know that $v(\rho_{2})=\lambda$ is a non zero constant. This implies that
$v(D_{0}(f)_{d-3})=S_{d-3-d_{2}}\subset B({\mathcal{A}},\rho).$
Any element $D^{\prime}\in D_{0}(f)_{d-3}$ can be written by Theorem 1.4 as a sum
$D^{\prime}=\sum A_{p}\tilde{D}_{p},$
where $A_{p}\in S_{m_{p}-3}$ and $p$ runs through all the multiple points of ${\mathcal{A}}$ satisfying $m_{p}\geq 3$. Hence
$v(D^{\prime})=\sum A_{p}v(\tilde{D}_{p})=\sum A_{p}g_{p}.$
Since in this sum
$0\leq\deg g_{p}=d_{1}+1-m_{p}\leq\frac{d+1}{2}-m_{p},$
it follows that the points $p$ in the above sum satisfy (1.5). This shows that
$S_{d-3-d_{2}}\subset I(\rho),$
and hence the zero set of the ideal $I(\rho)$ is empty in $\mathbb{P}^{2}$. Conversely, if the zero set of the ideal $I(\rho)$ is empty in $\mathbb{P}^{2}$, then the same holds for the ideal $B({\mathcal{A}},\rho)$, since by definition
$I(\rho)\subset B({\mathcal{A}},\rho).$
But the ideal $B({\mathcal{A}},\rho)$ being saturated, see [15, Theorem 5.1], this is possible if and only if $B({\mathcal{A}},\rho)=S$. Using (5.3), we conclude that ${\mathcal{A}}$ is free, and this completes the proof of Theorem 1.8.
5.3. Proof of Proposition 1.9
We choose $\rho=\tilde{D}_{q}$ as the minimal degree derivation and the coordinates such that $q=(1:0:0)$ and $p=(0:0:1)$. Then $\Delta(\tilde{D}_{p})$ is equal to the determinant of the $3\times 3$ matrix having as first row $x,y,z$, as second row $(f_{2q},0,0)$ and the third row $(0,0,f_{2p})$, and hence, up to a sign, one has
$\Delta(\tilde{D}_{p})=yf_{2p}f_{2q}.$
Case 1: The line $L(p,q):y=0$ is in ${\mathcal{A}}$. Then one can write $f_{1p}=yf_{1p}^{\prime}$, such that all the lines defined by $f_{1p}^{\prime}=0$ are in ${\mathcal{A}}$ and do not contained $q$. It follows that
$f_{2q}=f_{1p}^{\prime}h,$
where $h=0$ is the union of lines containing neither $p$ nor $q$. It follows that $h=0$ is an equation for the arrangement ${\mathcal{B}}_{p}$ defined above, and hence $h=h_{p}$ up to a constant factor. On the other hand we have
$g_{p}=\frac{\Delta(\tilde{D}_{p})}{f}=\frac{yf_{2p}f_{2q}}{f_{1p}f_{2p}}=\frac{yf_{2p}f_{1p}^{\prime}h}{f_{1p}f_{2p}}=h,$
which proves our claim in this case.
Case 2: The line $L(p,q):y=0$ is not in ${\mathcal{A}}$. Then one can write as above $f_{2q}=f_{1p}h$, where where $h=0$ is the union of lines containing neither $p$ nor $q$. It follows that $yh=0$ is an equation for the arrangement ${\mathcal{B}}_{p}$ defined above, and hence $h_{p}=yh$ up to a constant factor. On the other hand we have
$g_{p}=\frac{\Delta(\tilde{D}_{p})}{f}=\frac{yf_{2p}f_{2q}}{f_{1p}f_{2p}}=\frac{yf_{2p}f_{1p}h}{f_{1p}f_{2p}}=yh,$
which proves our claim in this case and completes thus the proof of Proposition 1.9.
Remark 5.4.
Note that if $m\leq d/2$ and
$\tau({\mathcal{A}})=(d-1)^{2}-m(d-m-1),$
then the arrangement ${\mathcal{A}}$ is free with exponents $(m,d-m-1)$. Indeed, using [17, 6] it follows that $d_{1}\leq m$. Theorem 1.8 $(2)$, implies that $d_{1}\geq m-1$, and the equality $d_{1}=m-1$ is not possible. Indeed, again by Theorem 1.8 $(2)$, $d_{1}=m-1$ implies that
$\tau({\mathcal{A}})=(d-1)^{2}-(m-1)(d-m)>(d-1)^{2}-m(d-m-1),$
a contradiction. Therefore, the simplest case where the freeness cannot be decided easily by the combinatorics seems to be the case
$d_{1}=m+1\text{ and }2m+3\leq d.$
For instance, the free arrangement ${\mathcal{A}}$ with exponents $(m+1,m+1)$ and the nearly free arrangement ${\mathcal{B}}$ with exponents $(m,m+3,m+3)$ satisfy $\tau({\mathcal{A}})=\tau({\mathcal{B}})$. We do not know whether such a pair of arrangements ${\mathcal{A}},{\mathcal{B}}$ exists, such that ${\mathcal{A}}$ and ${\mathcal{B}}$ have the same combinatorics.
6. Derivations of Euler type, the monomial and the full monomial line arrangements
We say that a derivation $D_{\rho}$ as in (1.2) is of Euler type if
$a=xA,\ b=yB\text{ and }c=zC$
for some polynomials $A,B,C\in S$. Assume that
(6.1) $p=(1:0:0),\ p^{\prime}=(0:1:0)\text{ and }p^{\prime\prime}=(0:0:1)$
are multiple points in the line arrangement ${\mathcal{A}}$, with respective multiplicities $m\geq m^{\prime}\geq m^{\prime\prime}\geq 3$. To simplify notation, we set
$\Delta=\det M_{p},\ \Delta^{\prime}=\det M_{p^{\prime}}\text{ and }\Delta^{\prime\prime}=\det M_{p^{\prime\prime}}$
where the matrix $M_{p}$ is as defined in section 4, and
$g=v(\tilde{D}_{p}),\ g^{\prime}=v(\tilde{D}_{p^{\prime}})\text{ and }g^{\prime\prime}=v(\tilde{D}_{p^{\prime\prime}}).$
Lemma 6.1.
With the above notation one has the following.
  1. (1)
    If the minimal degree derivation $\rho$ of the arrangement ${\mathcal{A}}$ is given by $(a,b,c)$, then
    $\Delta=yc-zb,\ \Delta^{\prime}=za-xc\text{ and }\Delta^{\prime\prime}=xb-ya.$
    For any other multiple point $q=(u:v:w)\in{\mathcal{A}}$ one has
    $\Delta_{q}=u\Delta+v\Delta^{\prime}+w\Delta^{\prime\prime},$
    where $\Delta_{q}=\Delta(\tilde{D}_{q})$.
  2. (2)
    If the minimal degree derivation $\rho$ of the arrangement ${\mathcal{A}}$ is of Euler type, say $\rho=(xA,yB,zC)$, then
    $\Delta=yz(C-B),\ \Delta^{\prime}=xz(A-C)\text{ and }\Delta^{\prime\prime}=xy(B-A).$
  3. (3)
    If the 3 lines determined by $p,p^{\prime}$ and $p^{\prime\prime}$ are in the arrangement ${\mathcal{A}}$, then any derivation $\rho$ of the arrangement ${\mathcal{A}}$ is of Euler type.
Proof.
The claims $(1)$ and $(2)$ are obvious by direct computation. To prove $(3)$, assume that $\rho=(a,b,c)$. Our assumption implies that $f=xyzh$, where $h$ is not divisible by $x,y$ or $z$. It follows that $f_{x}$ is not divisible by $x$, but both $f_{y}$ and $f_{z}$ are divisible by $x$. Hence the relation
$af_{x}+bf_{y}+cf_{z}=0$
implies that $a$ is divisible by $x$. In the same way we get that $b$ (resp. $c$) is divisible by $y$ (resp. by $z$). Hence $\rho$ is an Euler derivation. ∎
Remark 6.2.
The notion of derivation of Euler type depends on the choice of coordinates $(x:y:z)$ on $\mathbb{P}^{2}$, in other words on the choice of the three multiple points $p,p^{\prime}$ and $p^{\prime\prime}$. Consider the line arrangement
${\mathcal{C}}:f=(x^{3}+y^{3}+z^{3})^{3}-27x^{3}y^{3}z^{3}=0,$
which is a subarrangement of the Hessian line arrangement, see [5, Example 8.6]. A direct computation shows that ${\mathcal{C}}$ is a 4-syzygy curve with exponents $(4,6,6,6)$ such that the minimal degree derivation is of Euler type with
$A=-(y^{3}-z^{3}),\ B=(x^{3}-z^{3})\text{ and }C=-(x^{3}-y^{3}).$
However, the points $p,p^{\prime}$ and $p^{\prime\prime}$ as defined in (6.1) are not situated on ${\mathcal{C}}$. It would be more useful in this case to choose
$p=(1:e:0),\ p^{\prime}=(0:1:e)\text{ and }p^{\prime\prime}=(e^{2}:0:1),$
where $e=\exp(2\pi i/6)$. Indeed, these new 3 points are triple points of the arrangement ${\mathcal{C}}$, since they are solutions of the equations
$x^{3}+y^{3}+z^{3}=xyz=0.$
The 3 lines determined by these points are in ${\mathcal{C}}$, so one may choose their equations as a new coordinate system on $\mathbb{P}^{2}$, and in this new coordinate system $p,p^{\prime}$ and $p^{\prime\prime}$ are given by (6.1).
Example 6.3.
Consider the full monomial arrangement
${\mathcal{F}}_{m}:xyz(x^{m-2}-y^{m-2})(x^{m-2}-y^{m-2})=0,$
which is free with exponents $(m-1,2m-3)$ for $m\geq 3$. This is a $(3,2m)$-multinet as explained for instance in [5, Example 6.12]. The minimal degree derivation $\rho$ is given by
$\rho=(xA,yB,zC),$
where
$A=mx^{m-2}-(2m-3)y^{m-2}-(2m-3)z^{m-2},\ B=-(2m-3)x^{m-2}+my^{m-2}-(2m-3)z^{m-2}$
and
$C=-(2m-3)x^{m-2}-(2m-3)y^{m-2}+my^{m-2},$
see [5, Example 8.6]. Hence $\rho$ is of Euler type as predicted by Lemma 6.1 $(3)$.
Remark 6.4.
Note that the minimal degree derivation $\rho$ of the arrangement ${\mathcal{A}}$ can be of Euler type even when the condition in $(3)$ above does not hold. For instance, in the monomial arrangement
${\mathcal{A}}_{m}:f=(x^{m}-y^{m})(y^{m}-z^{m})(x^{m}-z^{m})=0,$
the minimal degree derivation $\rho$ for $m\geq 4$ is given by
$\rho=(x(x^{m}-2y^{m}-2z^{m}),y(y^{m}-2x^{m}-2z^{m}),z(z^{m}-2x^{m}-2y^{m})),$
see [5, Example 8.6]. Hence $\rho$ is of Euler type, even if the lines $x=0$, $y=0$ and $z=0$ are not in ${\mathcal{A}}_{m}$. Note that the monomial arrangement ${\mathcal{A}}_{m}$, also called the Ceva arrangement, is a $(3,m)$-net, as explained for instance in [5, Example 6.11].
Theorem 6.5.
With the above notation, assume that $m=m^{\prime}=m^{\prime\prime}\geq 4$ is the maximal multiplicity of points in ${\mathcal{A}}$ and that ${\mathcal{A}}$ has a minimal degree derivation $\rho$ which is of Euler type and such that $d_{1}=\deg\rho=m+1$. If none of the lines $x=0$, $y=0$ and $z=0$ is in ${\mathcal{A}}$, then the arrangement ${\mathcal{A}}$ is projectively equivalent to the monomial arrangement ${\mathcal{A}}_{m}$.
Proof.
First notice that
$\deg\Delta=d_{1}+1=m+2\text{ and }\deg f_{1p}=m.$
Since $\Delta=yz(C-B)$ and $y,z$ are not factors of $f_{1p}$, it follows that $f_{1p}$ divides $C-B$. Since both polynomials have degree $m$, it follows that there is $k\in\mathbb{C}^{*}$ such that $C-B=kf_{1p}$. The other two equalities
$A-C=k^{\prime}f_{1p^{\prime}}\text{ and }B-A=k^{\prime\prime}f_{1p^{\prime\prime}}$
are proved in exactly the same way, where $k^{\prime},k^{\prime\prime}\in\mathbb{C}^{*}$. The equalities
$C-B=kf_{1p},\ A-C=k^{\prime}f_{1p^{\prime}}\text{ and }B-A=k^{\prime\prime}f_{1p^{\prime\prime}}$
imply that the arrangement ${\mathcal{A}}^{\prime}:f^{\prime}=0$ with $f^{\prime}=f_{1p}f_{1p^{\prime}}f_{1p^{\prime\prime}}$ is a $(3,m)$-net type arrangement, see for instance in [5, Theorem 6.5]. If we add to ${\mathcal{A}}^{\prime}$ the 3 lines given by $x=0$, $y=0$ and $z=0$, we get a supersolvable line arrangement ${\mathcal{B}}$ with 3 modular points, all of multiplicity $m+2$. Using [1, Theorem 1.2] or [18], we see that ${\mathcal{B}}$ is projectively equivalent to the full monomial arrangement
${\mathcal{F}}_{m+2}:xyz(x^{m}-y^{m})(y^{m}-z^{m})(x^{m}-z^{m})=0.$
It follows that ${\mathcal{A}}^{\prime}$ is obtained from the monomial arrangement ${\mathcal{A}}_{m}$ by an automorphism $\phi$ of $\mathbb{P}^{2}$ preserving the set $T=\{p,p^{\prime},p^{\prime\prime}\}$, that is $\phi(T)=T$. To complete the proof of Theorem 6.5 we have to show that ${\mathcal{A}}^{\prime}={\mathcal{A}}$. To do this, let $L$ be a line in ${\mathcal{A}}$ which is not in ${\mathcal{A}}^{\prime}$. Then $L$ cannot pass through any of the points $p,p^{\prime},p^{\prime\prime}$, since these points have already multiplicity $m$ in ${\mathcal{A}}^{\prime}$. We use then the following.
Lemma 6.6.
Consider the torus
${\mathcal{T}}=\{(x,y)\in\mathbb{C}^{2}\ :\ |x|=|y|=1\}.$
Then any line $\ell$ in $\mathbb{C}^{2}$ given by an equation
$ax+by+1=0$
with $ab\neq 0$ intersects the torus ${\mathcal{T}}$ in at most two points.
Proof.
To prove this Lemma, assume that $(u,v)\in{\mathcal{T}}\cap\ell.$ Then one has
$au+bv+1=0.$
Using the identification between complex numbers in $\mathbb{C}$ and vectors in $\mathbb{R}^{2}$, the above relation says that the vectors $V_{1}=au$, $V_{2}=bv$ and $V_{3}=1$ are the (oriented) sides of triangle $\mathbb{T}$ in $\mathbb{R}^{2}$. Note that the lengths of these sides are $|a|,|b|$ and $1$, and hence independent of $(u,v)$. Since the third oriented side $V_{3}$ is fixed, there are at most two possibilities for the triangle $\mathbb{T}$, once we fix an order of the sides up-to a cyclic permutation. For instance the order $(V_{1},V_{2},V_{3})$ can produce two values for the oriented angle $\widehat{(V_{1},V_{3})}$, call them ${\alpha}$ and ${\beta}$, satisfying ${\alpha}+{\beta}=2\pi$. Since the angle $\widehat{(V_{1},V_{3})}$ determines the angle $\widehat{(V_{2},V_{3})}$ in the triangle $\mathbb{T}$, and since there is an obvious identification
$u=\widehat{(V_{1},V_{3})}\text{ and }v=\widehat{(V_{2},V_{3})}$
this completes the proof of Lemma 6.6. ∎
Now we come back to the proof of Theorem 6.5. We can assume that ${\mathcal{A}}^{\prime}={\mathcal{A}}_{m}$. Then all the triple points of the arrangement ${\mathcal{A}}^{\prime}$ are in the affine open set $U_{z}:z=1$. The equation of the line $\ell=L\cap U_{z}$ can be taken to be of the form $ax+by+1=0$ with $ab\neq 0$. Using Lemma 6.6, we see that at most 2 triple points of ${\mathcal{A}}^{\prime}$ can be situated on the line $L$. It follows that the number
$r=|{\mathcal{A}}^{\prime}\cap L|$
of intersection points between $L$ and ${\mathcal{A}}^{\prime}$ satisfies
$r=(3m-3n_{3})+n_{3}\geq 3m-4,$
where $n_{3}$ denotes here the number of triple points in ${\mathcal{A}}^{\prime}$ situated on $L$. Now we apply [10, Theorem 1.4] and deduce that the minimal degree $d_{1}^{\prime}$ of a Jacobian derivation for the arrangement ${\mathcal{A}}^{\prime\prime}={\mathcal{A}}^{\prime}\cup L$ is $d_{1}+1=m+2$. Indeed, $d_{1}^{\prime}=d_{1}$ would imply
$r=d_{1}+1=m+2,$
a contradiction since
$23m-4>m+$ 2
for $m\geq 4$. By adding new lines to any line arrangement the minimal degree of a Jacobian derivation cannot decrease, see for instance [13]. It follows that there is no line $L\in{\mathcal{A}}$ which is not in ${\mathcal{A}}^{\prime}$, and this completes the proof of Theorem 6.5. ∎
Example 6.7.
Consider the monomial arrangement ${\mathcal{A}}_{m}$ from Remark 6.4. In this case one has
$\Delta=yz(C-B)=3yz(z^{m}-y^{m}),\ \Delta^{\prime}=xz(A-C)=3xz(x^{m}-z^{m})\text{ and }\Delta^{\prime\prime}=3xy(y^{m}-x^{m}).$
For the triple point $q=(1:1:1)$ we have
$\Delta_{q}=-3[xy(x^{m}-y^{m})+yz(y^{m}-z^{m})+zx(z^{m}-x^{m})].$
A simple computation shows that
$g_{q}=\frac{\Delta_{q}}{(x-y)(y-z)(z-x)}=3P_{m-1}(x,y,z),$
where $P_{m-1}$ is the sum of all monomials in $x,y,z$ of degree $m-1$, that is
$P_{m-1}(x,y,z)=\sum_{({\alpha},{\beta},\gamma):{\alpha}+{\beta}+\gamma=m-1}x^{{\alpha}}y^{{\beta}}z^{\gamma}.$
In particular, the ideal generated by $xy,yz,xz$ and $g_{p}$ has no zero in $\mathbb{P}^{2}$, which gives a new proof for the freeness of the arrangement ${\mathcal{A}}_{m}$.
Theorem 6.8.
With the above notation, assume that $m=m^{\prime}=m^{\prime\prime}\geq 3$ is the maximal multiplicity of points in ${\mathcal{A}}$ and that ${\mathcal{A}}$ has a minimal degree derivation $\rho$ such that $d_{1}=\deg\rho=m-1$. If all the lines $x=0$, $y=0$ and $z=0$ are in ${\mathcal{A}}$, then the arrangement ${\mathcal{A}}$ is projectively equivalent to the full monomial arrangement ${\mathcal{F}}_{m}$.
Proof.
First notice that
$\deg\Delta=d_{1}+1=m=\deg f_{1p}.$
Using Lemma 6.1 $(3)$, we know that $\rho$ is of Euler type, say $\rho=(xA,yB,zC).$ Since $\Delta=yz(C-B)$ and $y,z$ are factors of $f_{1p}$, it follows that $f_{1p}^{\prime}=f_{1p}/yz$ divides $C-B$. Since both polynomials have degree $m-2$, it follows that there is $k\in\mathbb{C}^{*}$ such that $C-B=kf_{1p}^{\prime}$. The other two equalities
$A-C=k^{\prime}f_{1p^{\prime}}^{\prime}\text{ and }B-A=k^{\prime\prime}f_{1p^{\prime\prime}}^{\prime}$
are proved in exactly the same way, where $k^{\prime},k^{\prime\prime}\in\mathbb{C}^{*}$, $f_{1p^{\prime}}^{\prime}=f_{1p^{\prime}}/xz$ and $f_{1p^{\prime\prime}}^{\prime}=f_{1p^{\prime\prime}}/xy$. The equalities
$C-B=kf_{1p}^{\prime},\ A-C=k^{\prime}f_{1p^{\prime}}^{\prime}\text{ and }B-A=k^{\prime\prime}f_{1p^{\prime\prime}}^{\prime}$
imply that the arrangement ${\mathcal{A}}^{\prime}:f^{\prime}=0$ with $f^{\prime}=f_{1p}^{\prime}f_{1p^{\prime}}^{\prime}f_{1p^{\prime\prime}}^{\prime}$ is a $(3,m-2)$-net type arrangement, see for instance in [5, Theorem 6.5]. If we add to ${\mathcal{A}}^{\prime}$ the 3 lines given by $x=0$, $y=0$ and $z=0$, we get a supersolvable line arrangement ${\mathcal{B}}$ with 3 modular points, all of multiplicity $m\geq 3$. Using [1, Theorem 1.2] or [18], we see that ${\mathcal{B}}$ is projectively equivalent to the full monomial arrangement
${\mathcal{F}}_{m}:xyz(x^{m-2}-y^{m-2})(x^{m-2}-y^{m-2})=0.$
To complete the proof, we have to show that ${\mathcal{A}}={\mathcal{B}}$, that is that there is no additional line $L\in{\mathcal{A}}$ which is not in ${\mathcal{B}}$. If we add such a line to ${\mathcal{B}}$, we get a new line arrangement ${\mathcal{B}}^{\prime}$. Let $d_{1}=m-1$ (resp. $d_{1}^{\prime}$) be the minimal degree of a Jacobian derivation for ${\mathcal{B}}$ (resp. for ${\mathcal{B}}^{\prime}$). Now we apply [10, Theorem 1.4] and deduce that $d_{1}^{\prime}=d_{1}$ would imply
$r=d_{1}+1=m,$
where $r=|L\cap{\mathcal{B}}|$. Let $n_{k}$ be the number of $k$-multiple points of ${\mathcal{B}}={\mathcal{F}}_{m}$ situated on the new line $L$, for $k=2,3$. Then
$r=(3m-3)-3n_{3}-2n_{2}+n_{3}+n_{2}=3m-3-n_{3}-n_{2}=m+(m-3+m-n_{2}-n_{3})>m$
since $n_{3}\leq 2$ by Lemma 6.6 and a simple computation shows that $n_{3}+n_{2}\leq 2$ for $3\leq m\leq 4$, and $n_{3}+n_{2}\leq 3$ for $m\geq 5$ and for $m-2$ not divisible by 6. In such a case, when $m\geq 5$, for instance the following 3 points are on a line $L\notin{\mathcal{B}}$:
$(0:1:1),\ (1:v:w)\text{ and }(1:-w:-v)$
where $u,v$ are $m-2$-th roots of unity such that $v^{2}\neq w^{2}$. When $m-2$ is divisible by 6, hence $m\geq 8$, then $n_{3}+n_{2}\leq 4$. Indeed, in this case one can have $n_{3}=n_{2}=2$ as the following example shows. For $m=8$ the following 4 points are on a line $L\notin{\mathcal{B}}$:
$(0:1:1),\ (1:1:0),\ (1:e:e^{2})\text{ and }(1:e^{5}:e^{4})$
where $e=\exp(2\pi i/6)$. It follows that $d_{1}^{\prime}>d_{1}$ and the proof is completed as in the case of Theorem 6.5 above. ∎
We have the following purely combinatorial characterization of the full monomial arrangement ${\mathcal{F}}_{m}$.
Corollary 6.9.
With the above notation, assume that $m=m^{\prime}=m^{\prime\prime}\geq 3$ is the maximal multiplicity of points in ${\mathcal{A}}$ and all the lines $x=0$, $y=0$ and $z=0$ connecting these 3 points are in ${\mathcal{A}}$. Assume moreover that ${\mathcal{A}}$ contains $33m-$ 3 lines and that
$\tau({\mathcal{A}})\geq 7m^{2}-19m+13.$
Then this inequality is in fact an equality and the line arrangement ${\mathcal{A}}$ is projectively equivalent to the full monomial arrangement ${\mathcal{F}}_{m}$.
Proof.
Using Theorem 6.8, the only thing we should show is that $d_{1}=m-1$. Using the upper bound $\tau_{max}(d,d_{1})$ for the total Tjurina number $\tau({\mathcal{A}})$ in terms of $d=|{\mathcal{A}}|=3m-3$ and the minimal degree $d_{1}=\deg\rho$ given in [17], it follows that
$d_{1}\leq m-1.$
On the other hand, Theorem 1.8 implies that since $m\leq d/2$, we have $d_{1}\geq m-1$. It follows that $d_{1}=m-1$ and
$\tau({\mathcal{A}})=\tau_{max}(3m,m-1)=7m^{2}-19m+13.$
To get a corollary to Theorem 6.5 similar to Corollary 6.9, one may proceed as follows. We have seen in the proofs of Theorems 6.5 and 6.8 that for a minimal degree derivation of Euler type
(6.2) $D_{\rho}=xA\partial_{x}+yB\partial_{y}+zC\partial_{z}$
it is often possible to compute (to some extent) the differences $B-A$, $C-A$ and $B-C$. With this in view, the following result might be useful.
Proposition 6.10.
Assume that the differences $A-B$ and $A-C$ are known. Then the curve $C:f=0$ admits the Euler type derivation $D_{\rho}$ in (6.2), that is $D_{\rho}\in D_{0}(f)$, if and only if the polynomial
$h=y(A-B)f_{y}+z(A-C)f_{z}$
is divisible by $f$.
Proof.
Suppose first that the derivation $D_{\rho}$ is in $D_{0}(f)$. Then the derivation $D=D_{\rho}-AE$ satisfies
$D(f)=-dAf,$
where $d=\deg(f)$. On the other hand, we also have
$D(f)=y(B-A)f_{y}+z(C-A)f_{z}=-h.$
This implies that $h$ is divisible by $f$. Conversely, if $h$ is divisible by $f$, we set
$A=\frac{h}{df},B=(B-A)+\frac{h}{df}\text{ and }C=(C-A)+\frac{h}{df}.$
A direct computation shows that the derivation $D_{\rho}$ determined in this way belongs to $D_{0}(f)$. ∎
Example 6.11.
For the monomial arrangement ${\mathcal{A}}_{m}$ we have seen in the proof of Theorem 6.5 that one has
$A-C=k^{\prime}f_{1p^{\prime}}=k^{\prime}(x^{m}-z^{m})\text{ and }A-B=k^{\prime\prime}f_{1p^{\prime\prime}}=k^{\prime\prime}(x^{m}-y^{m}).$
In fact, the non zero constants can be chosen such that $k^{\prime}=k^{\prime\prime}=1$ and then
$h=y(A-B)f_{y}+z(A-C)f_{z}=my^{m}(x^{m}-y^{m})(x^{m}-z^{m})(x^{m}+z^{m}-2y^{m})+$
$+mz^{m}(x^{m}-z^{m})(x^{m}-y^{m})(2z^{m}-x^{m}-y^{m})=$
$=m(x^{m}-y^{m})(x^{m}-z^{m})(x^{m}y^{m}+y^{m}z^{m}-2y^{2m}+2z^{2m}-x^{m}z^{m}-y^{m}z^{m})=$
$=m(x^{m}-y^{m})(x^{m}-z^{m})(y^{m}-z^{m})(x^{m}-2y^{m}-2z^{m}).$
It follows that
$A=\frac{h}{df}=\frac{1}{3}(x^{m}-2y^{m}-2z^{m})$
$B=B-A+A=y^{m}-x^{m}+\frac{1}{3}(x^{m}-2y^{m}-2z^{m})=\frac{1}{3}(y^{m}-2x^{m}-2z^{m})$
and
$C=C-A+A=z^{m}-x^{m}+\frac{1}{3}(x^{m}-2y^{m}-2z^{m})=\frac{1}{3}(z^{m}-2x^{m}-2y^{m}).$
Up to the factor $\frac{1}{3}$, the derivation $D_{\rho}$ computed here is exactly the derivation given in Remark 6.4 above.
In fact, we have the following purely combinatorial characterization of the monomial arrangement ${\mathcal{A}}_{m}$.
Corollary 6.12.
With the above notation, assume that $m=m^{\prime}=m^{\prime\prime}\geq 3$ is the maximal multiplicity of points in ${\mathcal{A}}$ and that none of the lines $x=0$, $y=0$ and $z=0$ connecting these 3 points is in ${\mathcal{A}}$. Assume moreover that ${\mathcal{A}}$ contains $3m$ lines and that
$\tau({\mathcal{A}})\geq 7m^{2}-6m+3.$
Then this inequality is in fact an equality and the line arrangement ${\mathcal{A}}$ is projectively equivalent to the monomial arrangement ${\mathcal{A}}_{m}$.
Proof.
Using Theorem 6.5 it is enough to prove that $d_{1}=m+1$ and ${\mathcal{A}}$ has a minimal degree derivation $D_{\rho}$ which is of Euler type. The fact that $d_{1}=m+1$ follows exactly as in the proof of Corollary 6.9. To construct $\rho$ we use Proposition 6.10. If the minimal degree derivation $D_{\rho}$ is given by the formula (6.2), which $\deg A=\deg B=\deg C$, then it follows as in the proof of Theorem 6.5 that one has
(6.3) $C-B=f_{1p},\ A-C=f_{1p^{\prime}}\text{ and }B-A=f_{1p^{\prime\prime}}$
by choosing $k=k^{\prime}=k^{\prime\prime}=1$, and moreover $f=f_{1p}f_{1p^{\prime}}f_{1p^{\prime\prime}}$, since $\deg f=3m.$ Note that $f_{1p}\in\mathbb{C}[y,z]$, $f_{1p^{\prime}}\in\mathbb{C}[x,z]$ and $f_{1p^{\prime\prime}}\in\mathbb{C}[x,y]$. To apply Proposition 6.10, we compute the polynomial $h$ and we get
$h=-yf_{1p^{\prime\prime}}f_{1p^{\prime}}(f_{1p}f_{1p^{\prime\prime}})_{y}+zf_{1p^{\prime}}f_{1p^{\prime\prime}}(f_{1p}f_{1p^{\prime}})_{z}.$
To show that $h$ is divisible by $f$, we have to show that
$h^{\prime}=-y(f_{1p}f_{1p^{\prime\prime}})_{y}+z(f_{1p}f_{1p^{\prime}})_{z}$
is divisible by $f_{1p}$. Note that the equalities above in (6.3) imply that
$f_{1p}+f_{1p^{\prime}}+f_{1p^{\prime\prime}}=0$
and hence $f_{1p^{\prime\prime}}=-(f_{1p}+f_{1p^{\prime}})$. It follows that
$h^{\prime}=y(f_{1p}^{2})_{y}+y(f_{1p}f_{1p^{\prime}})_{y}+z(f_{1p}f_{1p^{\prime}})_{z}=$
$2yf_{1p}(f_{1p})_{y}+2mf_{1p}f_{1p^{\prime}},$
which is clearly divisible by $f_{1p}$. It follows that ${\mathcal{A}}$ has a minimal degree derivation $D_{\rho}$ which is of Euler type which coefficients $A,B,C$ are given by the formulas in Proposition 6.10 where $A-B=-f_{1p^{\prime\prime}}$ and $A-C=f_{1p^{\prime}}$.
Example 6.13.
If the condition
$\tau({\mathcal{A}})\geq 7m^{2}-6m+3$
is not satisfied, then $d_{1}=m+1$ fails, but the minimal degree derivation may still be of Euler type, as the following example shows. Consider the arrangement
$f=(x^{4}-y^{4})(y^{4}-z^{4})(x^{3}+x^{2}z+xz^{2}+z^{3})(x-2z)$
obtained from ${\mathcal{A}}_{4}$ by replacing the line $x-z=0$ by the line $x-2z=0$. Then ${\mathcal{A}}$ is a 4-syzygy curve with exponents $(6,7,7,8)$ and we have
$\tau({\mathcal{A}})=87<7\times 16-6\times 4+3=91.$
Moreover, the minimal degree derivation is of Euler type with $d_{1}=6>m+1$ and
$A=4x^{5}-8xy^{4}-9x^{4}z+16y^{4}z-x^{3}z^{2}-x^{2}z^{3}-9xz^{4}+16z^{5}$
$B=-8x^{5}+4xy^{4}+15x^{4}z-8y^{4}z-x^{3}z^{2}-x^{2}z^{3}-9xz^{4}+16z^{5}$
and
$C=-8x^{5}-8xy^{4}+15x^{4}z+16y^{4}z-x^{3}z^{2}-x^{2}z^{3}+3xz^{4}-8z^{5}.$
In this case one has
$A-B=12f_{1p^{\prime\prime}}(x-2z),\ A-C=12f_{1p^{\prime}}(x-z)\text{ and }B-C=12f_{1p}(x-2z).$
One can obtain a necessary condition for a line arrangement ${\mathcal{A}}$ to admit a minimal degree Euler type derivation as follows.
Proposition 6.14.
If the line arrangement ${\mathcal{A}}$ admits a derivation of Euler type
$\rho=(a,b,c)=(xA,yB,zC)$
of minimal degree $d_{1}$, then the number of multiple points of ${\mathcal{A}}$ outside the fundamental triangle $xyz=0$ is bounded by $(d_{1}-1)^{2}$.
Proof.
If $p$ is a multiple point of ${\mathcal{A}}$, it follows from the proof of [24, Proposition 3.6], see also Proposition 8.1 $(2)$ below, that the polynomials
$ya-xb=xy(A-B),\ za-xc=xz(A-C)\text{ and }zb-yc=yz(B-C)$
vanish at $p$. It follows that all the multiple points of ${\mathcal{A}}$ outside the union $xyz=0$ of coordinate lines are contained in the zero set
$Z:A-B=A-C=0.$
If $A-B$ and $A-C$ have a common factor $F$, then $F$ divides $B-C$, and hence the Bourbaki ideal $B({\mathcal{A}},\rho)$ is contained in the principal ideal $(F)$. This is a contradiction, since the zero set of $B({\mathcal{A}},\rho)$ is 0-dimensional, see [15, Theorem 5.1]. It follows that $Z$ is 0-dimensional complete intersection, and hence the number of multiple points of ${\mathcal{A}}$ outside the fundamental triangle $xyz=0$ is bounded by
$\deg(A-B)\deg(A-C)=(d_{1}-1)^{2}.$
Remark 6.15.
(i) The upper bound in Proposition 6.14 is sharp for the monomial arrangement ${\mathcal{A}}_{m}$ and for the full monomial arrangement ${\mathcal{F}}_{m}$.
(ii) The upper bound in Proposition 6.14 may be compared with the fact that the total number $M$ of multiple points of a free line arrangement ${\mathcal{A}}$ having a derivation $\rho=(a,b,c)$ of degree $e\geq 2$ and which is regular, that is $a,b,c$ is a regular sequence in $S$, satisfies
$M\leq e^{2}+e+1,$
see [24, Proposition 3.1]. The existence of such regular derivation with $e=d_{1}$ or $e=d_{2}$ is proven in [24, Theorem 2.2], where $(d_{1},d_{2})$ are the exponents of ${\mathcal{A}}$. The minimal degree Euler type derivation $\rho=(xA,yB,zC)$ considered in Proposition 6.14 is regular if and only if the system
$xA=yB=zC=0$
has no solution in $\mathbb{P}^{2}$.
7. Derivations of Euler type and addition-deletion of lines
The Euler type derivations behave well when we add certain lines. For instance we have the following result.
Proposition 7.1.
Assume that the degree $d$ curve $C:f=0$ has a minimal degree derivation
$\rho=(xA,yB,zC)$
of Euler type. Then the following hold.
  1. (1)
    If $x=0$ is not an irreducible component of the curve $C$, then the curve $C^{x}:f^{x}=xf=0$ has $\rho^{x}=(xA^{x},yB^{x},zC^{x})$ as a minimal degree derivation of Euler type, where
    $A^{x}=dA,\ B^{x}=(d+1)B-A\text{ and }C^{x}=(d+1)C-A.$
  2. (2)
    If $y=0$ is not an irreducible component of the curve $C$, then the curve $C^{y}:f^{y}=yf=0$ has $\rho^{y}=(xA^{y},yB^{y},zC^{y})$ as a minimal degree derivation of Euler type, where
    $A^{y}=(d+1)A-B,\ B^{y}=dB\text{ and }C^{y}=(d+1)C-B.$
  3. (3)
    If $z=0$ is not an irreducible component of the curve $C$, then the curve $C^{z}:f^{z}=zf=0$ has $\rho^{z}=(xA^{z},yB^{z},zC^{z})$ as a minimal degree derivation of Euler type, where
    $A^{z}=(d+1)A-C,\ B^{z}=(d+1)B-C\text{ and }C^{z}=dC.$
Proof.
The fact that $\rho^{x}$ is in $D_{0}(f^{x})$ can be checked by a direct obvious computation. If $d_{1}^{x}$ is the minimal degree of a derivation in $D_{0}(f^{x})$, then we know that $d_{1}^{x}\geq d_{1}=\deg\rho$. Since
$\deg\rho^{x}=\deg\rho=d_{1}$
it follows that $d_{1}^{x}=d_{1}$, unless $\rho^{x}=0$. But this case cannot occur, since $\rho^{x}=0$ yields $A=B=C=0$, which is impossible, as we have supposed $\rho\neq 0$. The other claims follows by replacing $x$ by $y$ and respectively by $z$. ∎
Example 7.2.
If we start with the monomial arrangement ${\mathcal{A}}_{m}$ for $m\geq 4$, we have
$A=x^{m}-2y^{m}-2z^{m},\ B=y^{m}-2x^{m}-2z^{m}\text{ and }C=z^{m}-2x^{m}-2y^{m},$
see Remark 6.4. The line arrangement
${\mathcal{A}}_{m}^{x}:f^{x}=x(x^{m}-y^{m})(y^{m}-z^{m})(x^{m}-z^{m})=0,$
obtained by adding the line $x=0$ has a minimal degree derivation $\rho^{x}$ of Euler type given by
$A^{x}=3mA,\ B^{x}=(3m+1)B-A\text{ and }C^{x}=(3m+1)C-A.$
Moreover, the line arrangement
${\mathcal{A}}_{m}^{xy}:f^{xy}=xy(x^{m}-y^{m})(y^{m}-z^{m})(x^{m}-z^{m})=0,$
obtained by adding the line $y=0$ to ${\mathcal{A}}_{m}^{x}$ has a minimal degree derivation $\rho^{xy}$ of Euler type given by
$A^{xy}=(3m+2)A^{x}-B^{x}=(3m+1)((3m+1)A-B),$
$B^{xy}=(3m+1)B^{x}=(3m+1)((3m+1)B-A),$
and
$C^{xy}=(3m+2)C^{x}-B^{x}=(3m+1)((3m+2)C-A-B).$
Finally, the line arrangement
${\mathcal{A}}_{m}^{xyz}:f^{xyz}=xyz(x^{m}-y^{m})(y^{m}-z^{m})(x^{m}-z^{m})=0,$
obtained by adding the line $z=0$ to ${\mathcal{A}}_{m}^{xy}$ has a minimal degree derivation $\rho^{xyz}$ of Euler type given by
$A^{xyz}=(3m+3)A^{xy}-C^{xy}=(3m+1)(3m+2)((3m+2)A-B-C),$
$B^{xyz}=(3m+1)B^{xy}-C^{xy}=(3m+1)(3m+2)((3m+2)B-A-C)$
and
$C^{xyz}=(3m+2)C^{xy}=(3m+1)(3m+2)((3m+2)C-A-B).$
Note that ${\mathcal{A}}_{m}^{xyz}$ is just the full monomial arrangement ${\mathcal{F}}_{m+2}$ and the minimal degree derivation $\rho^{xyz}$ constructed here coincides with the derivation $\rho$ given in Example 6.3, up to a constant factor $3(3m+1)(3m+2)$ and replacing $m$ by $m+2$.
It is easy to see that the line arrangements ${\mathcal{A}}_{m}^{x}$ and ${\mathcal{A}}_{m}^{xy}$ are both free, either by computing the total Tjurina number and using Remark 5.4, or by applying [10, Theorem 1.4].
Proposition 7.3.
Assume that the curve $C:f=0$ has a minimal degree derivation
${\rho}=(xA,yB,zC)$
of Euler type and that the line $L$ is an irreducible component of $C$. Then the curve $C^{\prime}:f^{\prime}=0$ obtained by deleting the line $L$ from $C$ has a derivation
$\rho^{\prime}=(xA^{\prime},yB^{\prime},zC^{\prime})$
of Euler type such that $A^{\prime}-B^{\prime}=A-B$ and $A^{\prime}-C^{\prime}=A-C$.
Proof.
Let $\ell={\alpha}x+{\beta}y+\gamma z$ and hence $f=\ell f^{\prime}$. It follows that
$f_{y}={\beta}f^{\prime}+\ell f_{y}^{\prime}\text{ and }f_{z}=\gamma f^{\prime}+\ell f_{z}^{\prime}.$
Now we apply Proposition 6.10 and compute
$h^{\prime}=y(A-B)f_{y}^{\prime}+z(A-C)f_{z}^{\prime}=\frac{y(A-B)(f_{y}-{\beta}f^{\prime})+z(A-C)(f_{z}-\gamma f^{\prime})}{\ell}=$
$=\frac{y(A-B)f_{y}+z(A-C)f_{z}}{\ell}-\frac{({\beta}y(A-B)+\gamma z(A-C))f^{\prime}}{\ell}.$
By Proposition 6.10 we know that
$h=y(A-B)f_{y}+z(A-C)f_{z}$
is divisible by $f$, and hence
$\frac{y(A-B)f_{y}+z(A-C)f_{z}}{\ell}$
is a polynomial divisible by $f^{\prime}$. It follows that
$\frac{({\beta}y(A-B)+\gamma z(A-C))f^{\prime}}{\ell}$
is also a polynomial, which is divisible by $f^{\prime}$, since $f^{\prime}$ and $\ell$ are relatively prime. ∎
Remark 7.4.
Let $d_{1}=\deg\rho=\deg\rho^{\prime}$ be the minimal degree of a derivation in $D_{0}(f)$. Let $d_{1}^{\prime}\leq d_{2}^{\prime}$ be the two smallest exponents of $C^{\prime}$. Then it is clear that
$d_{1}-1\leq d_{1}^{\prime}\leq d_{1}.$
Moreover, if $\rho^{\prime}$ is a primitive derivation, that is if $A^{\prime},B^{\prime}$ and $C^{\prime}$ have no common factor, and in addition $d_{2}^{\prime}>d_{1}$, then $d_{1}^{\prime}=d_{1}$, that is $\rho^{\prime}$ is a minimal degree derivation in $D_{0}(f^{\prime})$. In Example 6.13, we start with the line arrangement ${\mathcal{A}}_{4}$ and delete the line $L:x-z=0$ to get a new arrangement, call it ${\mathcal{A}}^{\prime}:f^{\prime}=0$. Then $|L\cap{\mathcal{A}}^{\prime}|=5$ and using [10, Theorem 1.3 (3)], we see that ${\mathcal{A}}^{\prime}$ is a plus-one arrangement with exponents $(5,6,6)$, in fact a nearly free arrangement. Hence in this case
$d_{1}=d_{1}^{\prime}=5$
and hence the derivation $D_{\rho^{\prime}}\in D_{0}(f^{\prime})$, identified to $\rho^{\prime}$ and constructed as in Proposition 7.1 has minimal degree in this case. Moreover, the line arrangement ${\mathcal{A}}$ from Example 6.13 has a minimal degree Euler type derivation of degree $6>5=\deg D_{\rho^{\prime}}$, which is up-to a constant factor just the derivation
$(x-2z)D_{\rho^{\prime}}-\frac{D_{\rho^{\prime}}(x-2z)}{12}E,$
coming from the derivation $D_{\rho^{\prime}}$ as explained in [13, Theorem 5.1]. This explains to some extent why Proposition 7.1 holds only for the special lines $x=0$, $y=0$ or $z=0$.
8. Proof of Theorems 1.10 and 1.14 and of Corollaries 1.11 and 1.13
We start with the following preliminary result.
Proposition 8.1.
Let ${\mathcal{A}}:f=0$ be a line arrangement and $\rho=(a,b,c)\in D_{0}(f)$. Then the following holds.
  1. (1)
    If $L:\ell={\alpha}x+{\beta}y+\gamma z=0$ is one of the lines in ${\mathcal{A}}$, then
    $\sigma={\alpha}a+{\beta}b+\gamma c$
    is divisible by $\ell$.
  2. (2)
    If $p\in{\mathcal{A}}$ is a point of multiplicity at least $2$, then the vector $\rho(p)=(a(p),b(p),c(p))$ is either zero, or a multiple of $p$.
  3. (3)
    Let $q\in{\mathcal{A}}$ be a point of multiplicity at least $2$ and such that $p\neq q$ and the line determined by $p$ and $q$ is not in ${\mathcal{A}}$. Then $g_{p}(q)=0$.
Note that the claims $(1)$ and $(2)$ are well known, see for instance [24, Proposition 3.6] and its proof.
Proof.
To prove $(1)$, we note that if we set $f=\ell f^{\prime}$, then $D_{\rho}(f)$ is a sum of terms divisible by $\ell$ plus the sum $f^{\prime}\sigma$. It follows that $\sigma$ is divisible by $\ell$. If we denote by $L^{0}$ the hyperplane in $\mathbb{C}^{3}$ given by $\ell=0$, then the claim $(1)$ says exactly that $D_{\rho}$, regarded as a vector field on $\mathbb{C}^{3}$, is everywhere tangent to $L^{0}$.
To prove $(2)$, note that if $p\in L_{1}\cap L_{2}$, then the tangent vector $D_{\rho}(p)$ is in the line $L_{1}^{0}\cap L_{2}^{0}=\mathbb{C}\cdot p$. To prove $(3)$, we recall that
$\Delta_{p}=g_{p}f_{1p},$
see (5.4). If we evaluate this equality at the point $q$ we have $f_{1p}(q)\neq 0$, since the line determined by $p$ and $q$ is not in ${\mathcal{A}}$. Moreover we have $\Delta_{p}(q)=0$. Indeed, the first two rows in the matrix used to compute $\Delta_{p}(q)$ are $v_{1}=(u,v,w)$ and $v_{2}=(a(q),b(q),c(q))$, where we assume that $q=(u:v:w)$. The vector $v_{2}$ is a multiple of the vector $v_{1}$ by the claim $(2)$, and hence this determinant is zero. ∎
8.2. Proof of Theorem 1.10
The first claims follow directly from Proposition 8.1. Note that the inequality $d_{1}<d-m_{p}$ implies that $\tilde{D}_{p}$ is not a constant multiple of the minimal degree derivation $D_{\rho}$, and hence $g_{p}\neq 0$. To prove the inequalities, one uses Bezout’s Theorem and has to notice that a point $q$ which is not connected to $p$ by lines in ${\mathcal{A}}$ is situated in fact in the arrangement ${\mathcal{A}}_{2}:f_{2p}=0$ with $\deg f_{2p}=d-m_{p}$. Moreover, $q$ has the same multiplicity in ${\mathcal{A}}$ and in ${\mathcal{A}}_{2}$. If the line $L$ is in ${\mathcal{A}}$, it is clear that it is in fact a line in ${\mathcal{A}}_{2}$ as well.
Example 8.3.
(i) Consider the Hessian line arrangement
${\mathcal{H}}:xyz\left((x^{3}+y^{3}+z^{3})^{3}-27x^{3}y^{3}z^{3}\right)=0$
which is free and it has $d_{1}=4$, see [5, Example (8.6) (iii)]. Note that ${\mathcal{H}}$ is the union of 4 line arrangements ${\mathcal{A}}_{1},{\mathcal{A}}_{2},{\mathcal{A}}_{3}$ and ${\mathcal{A}}_{4}$, each one consisting of one triangle. More precisely,
${\mathcal{A}}_{j}:x^{3}+y^{3}+z^{3}-3\epsilon^{j}xyz=0,$
for $j=1,2,3$ and $\epsilon=\exp(2\pi i/3)$, and ${\mathcal{A}}_{4}:xyz=0$. The point $p=(0:1:-1)$, has multiplicity $m_{p}=4$, and the 4 lines passing through $p$ are $x=0$, $x+y+z=0$, $\epsilon x+y+z=0$ and $\epsilon^{2}x+y+z=0$, hence exactly one line in each triangle ${\mathcal{A}}_{j}$. It follows that in each triangle ${\mathcal{A}}_{j}$ there is a double point of ${\mathcal{H}}$ not situated on these 4 lines. These points are given by
$(1:\epsilon:\epsilon),(1:\epsilon^{2}:\epsilon^{2}),(1:1:1)\text{ and }(1:0:0).$
They are situated on the line $L:y-z=0$ which is not in ${\mathcal{H}}$. Next we show that there are no 5 lines in ${\mathcal{H}}$, containing all the multiple points of ${\mathcal{H}}$. Indeed, if such a set of lines exists, there is a triangle ${\mathcal{A}}_{k}$ containing only one line $L^{\prime}$ in this set. It follows that the double point in this triangle ${\mathcal{A}}_{k}$ which is not on the line $L^{\prime}$, is not contained in the union of the 5 lines.
(ii) Consider the full monomial arrangement
${\mathcal{F}}_{3}:xyz(x-y)(y-z)(x-z)=0.$
Then $d_{1}=2$ and let $p$ be the point $(0:1:1)$ of multiplicity 2. The points not connected by lines in ${\mathcal{F}}_{3}$ to the point $p$ are the points $(1:0:1)$ and $(1:1:0)$ which are situated on the line
$L:x-y-z=0,$
which is not a line in ${\mathcal{F}}_{3}$. Note that in this case
$\sum_{q}m_{q}=2+2=6-2=d-m_{p},$
and hence this inequality is sharp.
(iii) Consider next the supersolvable arrangement
${\mathcal{A}}:xyz(x^{2}-y^{2})(y-z)=0.$
Then $d_{1}=2$ and choose $p=(-1:1:1)$, which has multiplicity $m=2$. The points not connected by lines in ${\mathcal{A}}$ to the point $p$ are the points $(0:1:0)$ and $(1:1:0)$ which are situated on the line
$L:z=0,$
which is a line in ${\mathcal{A}}$. Note that in this case
$\sum_{q}(m_{q}-1)=1+1<3=6-2-1=d-m_{p}-1.$
(iv) Finally, let ${\mathcal{B}}_{1}$ (resp. ${\mathcal{B}}_{2}$) be the arrangement consisting of $n_{1}\geq 2$ lines passing through a point $a_{1}\in\mathbb{P}^{2}$ (resp. $n_{2}$ lines passing through a point $a_{2}\in\mathbb{P}^{2}\setminus{\mathcal{B}}_{1}$) and such that $n_{1}\leq n_{2}$. Then the arrangement ${\mathcal{A}}={\mathcal{B}}_{1}\cup{\mathcal{B}}_{2}$ has exponents $(n_{1},n_{2},n_{1}+n_{2}-2)$, see [2, Proposition 1.22]. If we choose $p=a_{1}$, then the only multiple point not connected by lines in ${\mathcal{A}}$ to the point $p$ is the point $a_{2}$. Any line through $a_{2}$ can be chosen as the line $L$.
8.4. Proof of Corollary 1.11
Choose a multiple point $p\in{\mathcal{A}}$ such that $d_{1}<d-m_{p}$, which as above implies that $g_{p}\neq 0$. With the notation from the proof of Theorem 1.10, a curve containing all the multiple points of ${\mathcal{A}}$ has equation $g_{p}f_{1p}=0$. Then ${\mathcal{C}}$ is obtained from this curve by deleting some irreducible components if they contain only multiple points contained already in the other irreducible components of ${\mathcal{C}}$.
8.5. Proof of Corollary 1.13
If $d_{1}<d-m_{p}$, then the claim follows from Theorem 1.10 $(2)$. On the other hand, if $d_{1}=d-m_{p}$, the claim follows from Theorem 1.14 which is proved below. Note that the inequality $d_{1}>d-m_{p}$ is impossible, since one has
$d_{1}\leq d-m\leq d-m_{p},$
where $m$ denotes the maximal multiplicity of points in ${\mathcal{A}}$, see [7, Theorem 1.2].
8.6. Proof of Theorem 1.14
If $d_{1}=d-m$, then we have $N\leq d_{1}+1=d-(m-1)$, since if we delete $m-1$ of the lines passing through $p$, the remaining $d-(m-1)$ lines contain all the multiple points. The equality $m-1=d_{1}$ implies that the ${\mathcal{A}}$ is supersolvable and such a point $p$ is modular, as shown in Theorem 1.10 (1). Hence the $m=d_{1}+1$ lines passing through $p$ contain all the multiple points.
In case $(1)$ it is known that $d_{1}=d-2$, see for instance [22] and hence $N\leq d-1$ by the claim proved above. On the other hand $N=d-1$, since if we delete 2 lines from ${\mathcal{A}}$ their intersection point is not on the union of the remaining $d-2$ lines.
In case $(2)$ we first show that $d_{1}=d-3$. Using Corollary 1.2, we see that $1\leq\dim D_{0}(f)_{d-3}\leq 3$. Since the local derivation $\tilde{D}_{p}$ associated with a point $p$ of multiplicity 3 is primitive, it follows that even when $\dim D_{0}(f)_{d-3}=3$, there is no nozero derivation of degree $<d-3$. Hence $d_{1}=d-3$ and hence $N\leq d-2$. Moreover, 3 lines cannot be deleted and preserving the required property, and hence $N=d-2$.
In case $(3)$, using [7, Theorem1.2], it follows that either
$d_{1}=d-m\text{ or }m-1=d_{1}.$
Hence we are exactly with the hypothesis of our Theorem 1.14.
In case $(4)$ let $p$ be a modular point such that its multiplicity $m$ is the highest multiplicity of points in ${\mathcal{A}}$, see [14, Lemma 3.1]. Then one has
$d_{1}=\min\{m-1,d-m\},$
see [14, Proposition 3.2]. So again we are exactly with the hypothesis of our Theorem 1.14.
Remark 8.7.
(i) Even when the curve ${\mathcal{C}}$ from Corollary 1.11 is a union of lines, some of the lines may not be in ${\mathcal{A}}$, as the first arrangement in Example 8.3 shows. For the monomial arrangement ${\mathcal{A}}_{m}$, if we take $p=(1:0:0)$ and use the formulas in Remark 6.2, the equation for ${\mathcal{C}}$ is the following
$\Delta_{p}=yz(y^{m}-z^{m})=0,$
and hence ${\mathcal{C}}$ contains two lines not in ${\mathcal{A}}_{m}$.
(ii) One may define another numerical invariant, namely $N_{0}$ to be the minimal number of lines in ${\mathcal{A}}$, containing all the multiple points of ${\mathcal{A}}$. It is clear that
$N_{0}\geq N,$
and Example 8.3 shows that this inequality may be strict.
In many cases the equality $N=d_{1}+1$ holds in Conjecture 1.12, but not always, as the following example shows.
Example 8.8.
The free line arrangement ${\mathcal{A}}_{m}^{x}$ from Example 7.2, has $d_{1}=m+1$ as shown there, and $N=d_{1}=m+1$, since the union of the $m$ lines through $(1:0:0)$ and the line $x=0$ contains all the multiple points of ${\mathcal{A}}_{m}^{x}$.